### v4

# Extremum Seeking Tracking for Derivative-free Distributed Optimization

Abstract
In this paper, we deal with a network of agents that want to cooperatively minimize the sum of local cost functions depending on a common decision variable. We consider the challenging scenario in which objective functions are unknown and agents have only access to local measurements of their local functions. We propose a novel distributed algorithm that combines a recent gradient tracking policy with an extremum-seeking technique to estimate the global descent direction. The joint use of these two techniques results in a distributed optimization scheme that provides arbitrarily accurate solution estimates through the combination of Lyapunov and averaging analysis approaches with consensus theory. We perform numerical simulations in a personalized optimization framework to corroborate the theoretical results.
I Introduction
In recent years, distributed optimization over networks has become a key research topic, see, e.g., [1, 2, 3] for an overview, also in the setting with partially unknown cost function [4]. Examples include data analytics in machine learning [5] as well as automatic controller tuning [6, 7].
We organize the literature review in two blocks: collaborative/distributed extremum seeking and distributed zeroth-order/derivative-free optimization. Early references on distributed extremum for the so-called consensus optimization framework are [8] for a discrete-time setting and [9] for a continuous-time one. A proportional-integral extremum-seeking design technique is proposed in [10], while, in [11], the gradient is approximated through a real-time protocol. In [12], authors propose the use of the sliding mode to generate the dither signal at the base of the extremum seeking. More recently, in [13], a distributed stochastic extremum seeking scheme is proposed for a source localization problem, while in [14] a distributed scheme approximating a Newton method is proposed. As for constraint-coupled distributed optimization, in [15], a distributed extremum seeking control, based on evolutionary game theory, is designed for real-time resource allocation. In [16], instead, a distributed, continuous-time scheme based on sign-based consensus is designed. In [17], a Lie bracket technique and extremum-seeking are used for problems with linear constraints. In [18], resource allocation problems are addressed by an extremum-seeking scheme.
As for distributed zeroth-order/derivative-free optimization, authors in [19] develop a zeroth-order scheme based on a one-point estimator and a gradient tracking policy. The work [20] instead proposes a zeroth-order algorithm based on a two-point estimator with a distributed gradient descent strategy and another one based on an $n$-point estimator with a gradient tracking policy. Authors in [21] propose a continuous-time gradient-free approach emulating a distributed gradient algorithm for which optimal asymptotic convergence is guaranteed. In [22], a sampled version of [21] is proposed. In [23], a continuous-time distributed algorithm based on random gradient-free oracles is proposed for convex optimization problems. In [24], the randomized gradient-free oracles introduced in [23] are used to build a gradient-free distributed algorithm in directed networks. Randomized gradient-free algorithms are used also in [25], where sequential Gaussian smoothing is used for non-smooth distributed convex constrained optimization. In [26], gradient-free optimization is addressed with the additional constraint that each agent can only transmit quantized information. The work in [27] instead develops a “directed-distributed projected pseudo-gradient” descent method for directed graphs. Paper [28] combines the gradient-free strategy of [23] with a saddle-point algorithm. Authors in [29] address an online constrained optimization problem by relying on the Kiefer-Wolfowitz algorithm to approximate the gradients, and [30] combines the estimation of the gradient via a “simultaneous perturbation stochastic approximation” technique with the so-called matrix exponential learning optimization method. In [31], a randomized gradient-free method is combined with a state-of-the-art distributed gradient descent approach for directed networks. In [32], an overview of zeroth-order methods based on a Frank-Wolfe framework is provided. Authors in [33] propose a distributed random gradient-free protocol to solve constrained optimization problems by using projection techniques. All the cited papers but [30] prove stability of the proposed schemes by using arguments not based on Lyapunov theory. We also note that [30] proposes an algorithm neither based on gradient tracking nor using extremum-seeking.
This paper proposes a novel algorithm to solve a distributed optimization problem in which network agents can only evaluate their local cost function at a given point, but not its gradient. The proposed solution consists of a distributed protocol in which the (unavailable) local gradients are approximated through an extremum seeking scheme. The approximations of the gradients are used to feed a suitable tracking mechanism which, in turn, allows for steering the local solution estimates along approximations of the global descent direction. The distributed algorithm uses a further consensus action on the local solution estimates. The convergence of our scheme is proved through Lyapunov and averaging theories for discrete-time systems. It is worth mentioning that our scheme, together with the ones in [34, 35], is the only distributed extremum-seeking scheme proposed in discrete-time with the following distinctive features. The work [34] (i) does not address a consensus optimization problem and (ii) relies on consensus dynamics estimating the global cost, while ours estimates the global gradient. Instead, in [35], (i) the addressed consensus optimization problems have scalar decision variables and (ii) an extremum-seeking technique is combined with a distributed gradient algorithm, i.e., without a tracking mechanism. As for the literature on distributed zeroth-order/derivative-free methods, the closest work is [19] in which extremum seeking is not used and the global gradient is approximated via a randomized one-point policy. Consistent with the comparison table provided in [19], we highlight that there are no other distributed algorithms in the literature using 1-point gradient estimators. Indeed, although the distributed scheme in [36] uses 1-point gradient estimators, we remark that it is tailored for partition-based optimization, i.e., a simplified setup in which each local function depends only on the neighbors’ decision variables, see [3]. Also, our Lyapunov-based tools combined with averaging theory for discrete-time systems represent a distinctive feature in the algorithm analysis.
The paper unfolds as follows. Section II introduces the problem and the proposed algorithm. The main result is provided in Section III and proved in Section IV. In Section V, we test the proposed algorithm via numerical simulations.
Notation
Given $N$ vectors $x_{1},\dots,x_{N}\in\mathbb{R}^{n}$, we denote by $\mathtt{col}(x_{1},\dots,x_{N})$ their column stacking. Given $N$ scalars $d_{1},\dots,d_{N}$, we denote by $\mathtt{diag}(d_{1},\dots,d_{N})$ the diagonal matrix with $i$-th entry $d_{i}$. The Kronecker product is denoted by $\otimes$. The identity matrix in ${\mathbb{R}}^{n\times n}$ is $I_{n}$. The column vector of $N$ ones is denoted by $1_{N}$ and we define $\mathbf{1}:=1_{N}\otimes I_{n}$. Dimensions are omitted whenever they are clear from the context. Finally, for $r>0$, we let $\mathcal{B}_{r}=\{x\in{\mathbb{R}}^{n}\,:\,\left\|x\right\|\leq r\}$.
II Problem Formulation
We consider a network of $N$ agents communicating according to an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{V}=\{1,\dots,N\}$ is the set of agents, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathcal{A}=[a_{ij}]\in{\mathbb{R}}^{N\times N}$ is the weighted adjacency matrix. Agent $i$ and $j$ can exchange information only if $(i,j)\in\mathcal{E}$. Accordingly, it holds $a_{ij}\geq 0$ if $(i,j)\in\mathcal{E}$ and $a_{ij}=0$ otherwise. We denote as $\mathcal{N}_{i}=\{j\in\mathcal{V}\mid(i,j)\in\mathcal{E}\}$ the set of neighbors of agent $i$. Moreover, we also associate to the graph the Laplacian matrix $\mathcal{L}:=D-\mathcal{A}\in{\mathbb{R}}^{N\times N}$, where $D:=\mathtt{diag}(\mathtt{deg}_{1},\dots,\mathtt{deg}_{N})\in{\mathbb{R}}^{N\times N}$ is the so-called degree matrix in which $\mathtt{deg}_{i}:=\sum_{j\in\mathcal{N}_{i}}a_{ij}$ is the degree of agent $i$. The next assumption specifies the class of considered graphs.
Assumption 1
The graph $\mathcal{G}$ is connected and the adjacency matrix $\mathcal{A}\in^{N\times N}$ is symmetric. $\square$
In the proposed distributed setup, each agent $i$ is equipped with a sensor only providing measurements of the local cost function $f_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and aims at solving the problem
$\min_{w\in\mathbb{R}^{n}}\sum_{i=1}^{N}f_{i}(w).$ (1)
We enforce the following assumptions about the problem.
Assumption 2
For all $i\in\{1,\ldots,N\}$, the function $f_{i}$ is $\underline{L}$-strongly convex for some $\underline{L}>0$. $\square$
Assumption 3
Each cost $f_{i}$ is $\mathcal{C}^{3}$ and has $\overline{L}_{i}$-Lipschitz continuous gradients. We denote $\overline{L}=\max\{\overline{L}_{1},\dots,\overline{L}_{N}\}$. $\square$
Assumption 2 ensures the existence of a unique solution $x^{\star}\in\mathbb{R}^{n}$ to problem (1). Our aim is to iteratively find it via a distributed algorithm. Namely, given the iteration index $t\in\mathbb{N}$ and by denoting with $w_{i}^{t}\in{\mathbb{R}}^{n}$ the $i$-th agent’s estimate, at iteration $t$, of the solution to problem (1), our goal is to design a distributed protocol able to steer all these estimates to the minimizer $x^{\star}$. The peculiar challenge of this paper is that each agent $i$ can access neither the gradients $\nabla f_{i}(w_{i}^{t})$ (as in standard gradient-based methods) nor the cost functions in arbitrary points (as in standard zeroth-order methods). More in detail, we assume that agent $i$ can only use the single measurement $f_{i}(w_{i}^{t})$ properly combined with the so-called dither signal $d_{i}^{t}\in{\mathbb{R}}^{n}$ and the amplitude parameter $\delta>0$. The role of $d_{i}^{t}$ and $\delta$ will become clearer in the next section.
III Extremum Seeking Tracking:
Algorithm Introduction and Convergence
Being the gradients $\nabla f_{i}$ not available, we replace them with an estimation based on a proper elaboration of the local costs excited via suitable dithering signals $d_{i}^{t}$. The arising distributed method, termed Extremum Seeking Tracking, is described in Algorithm 1, from agent $i$ perspective. In Algorithm 1, $\gamma>0$ represents the step size while the parameter $\delta>0$ represents the amplitude of the dither signal $d_{i}^{t}$ defined as
$d_{i}^{t}=\mathtt{col}\bigg{(}\sin\bigg{(}\dfrac{2\pi t}{\tau_{i_{1}}}+\phi_{i_{1}}\bigg{)},\,\dots,\,\sin\bigg{(}\dfrac{2\pi t}{\tau_{i_{n}}}+\phi_{i_{n}}\bigg{)}\bigg{)},$ (2)
where $\tau_{i_{p}}\in\mathbb{N}$ and $\phi_{i_{p}}\in\mathbb{R}$ such that, given $p,q,r\in\{1,\ldots,n\}$, $p\neq q$, $q\neq r$, $p\neq r$, it holds
$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\left(\tfrac{2\pi t}{\tau_{i_{p}}}+\phi_{i_{p}}\right)\!=\!0$ (3a)
$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\left(\tfrac{2\pi t}{\tau_{i_{p}}}+\phi_{i_{p}}\right)\sin\left(\tfrac{2\pi t}{\tau_{i_{q}}}+\phi_{i_{q}}\right)\!=\!\dfrac{\tau_{\text{per}}}{2}$ (3b)
$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{r}}}\!+\!\phi_{i_{r}}\right)\!=\!0,$ (3c)
for $i\in\{1,\ldots,N\}$. Here, $\tau_{\text{per}}\in\mathbb{N}$ is the least common multiple of all periods $\tau_{ip}$. Sinusoidal dither functions are useful in practical applications to guarantee smooth inputs to the plant. However, other possibilities as, e.g., as square or triangular waves, are possibile [37].
initialization: $x_{i}^{0}\in\mathbb{R}^{n}$ and $z_{i}^{0}=0$
for $t=0,1,\dots$ do
$w_{i}^{t+1}$ $=w_{i}^{t}-\gamma\sum_{j\in\mathcal{N}_{i}}\!\!{\ell}_{ij}(w_{j}^{t}-\delta d_{j}^{t})-\gamma s_{i}^{t}+\delta(d_{i}^{t+1}-d_{i}^{t})$ (4a)
$s_{i}^{t+1}$ $=s_{i}^{t}-\!\gamma\!\sum_{j\in\mathcal{N}_{i}}\ell_{ij}s_{j}^{t}+\frac{2}{\delta}(f_{i}(w_{i}^{t+1})d_{i}^{t+1}-f_{i}(w_{i}^{t})d_{i}^{t})\!$ (4b)
end for
Algorithm 1 Extremum Seeking Tracking (agent $i$)
The local gradient estimates generated by $\frac{f_{i}(w_{i}^{t})d_{i}^{t}}{2\delta}$ are suitably interlaced with (i) the term $\sum_{j\in\mathcal{N}_{i}}{\ell}_{ij}(w_{j}^{t}-\delta d_{j}^{t})$ to force consensus among the quantities $w_{i}^{t}-\delta d_{i}^{t}$ and (ii) a tracking mechanism to reconstruct the (estimated) global gradient. In detail, the consensus step is performed by using the entries $\ell_{ij}$ the $(i,j)$-entry of the Laplacian matrix $\mathcal{L}$ associated to the graph $\mathcal{G}$, while the tracking mechanism is implemented by equipping each agent $i$ with an auxiliary variable $s_{i}^{t}\in{\mathbb{R}}^{n}$. In this algorithm, agents exchange with their neighbors the information $\mathtt{col}(w_{i}^{t}-\delta d_{i}^{t},s_{i}^{t})$ involving $2n$ components.
Remark 1
The main distinctive features of our method are as follows. First, errors in our method are given by third-order residuals as opposed to second-order ones in finite-difference methods. Second, gradient estimation is based on a single function query. This could be advantageous in scenarios in which multiple queries are expensive or even not allowed. Third and final, the estimation policy updates, and so the convergence guarantees, are purely deterministic. $\square$
The next theorem formalizes the convergence properties of Extremum Seeking Tracking. To this end, for all $i\in\{1,\ldots,N\}$ and given $r>0$, we define the set $\mathcal{D}_{i,r}\subset{\mathbb{R}}^{2n}$ as
$\mathcal{D}_{i,r}\!:=\!\left\{\mathtt{col}(w_{i},z_{i})\in{\mathbb{R}}^{2n}\!\mid\!\left\|w_{i}-x^{\star}\right\|\leq r,s_{i}=2f_{i}(w_{i})/\delta\right\}\!.$
Theorem 1
Consider (4) and let Assumptions 12, and 3 hold. Then, for any $r,\bar{\rho}>0$, there exist $\gamma^{\star},\delta^{\star},k_{1}>0$, $\epsilon\in(0,\bar{\rho}/2)$, and $k_{2}\geq(\bar{\rho}/2-\epsilon)$ such that, for any $\gamma\in(0,\gamma^{\star})$ and $\delta\in(0,\delta^{\star})$, if $\mathtt{col}(w_{i}^{0},s_{i}^{0})\in\mathcal{D}_{i,r}$ for all $i\in\{1,\ldots,N\}$, the trajectories of (4) are bounded and for all $i\in\{1,\ldots,N\}$
$\left\|w_{i}^{t}-x^{\star}\right\|\leq\bar{\rho},$ (5)
for all $t\geq t^{\star}$, where $t^{\star}\!:=\!-\frac{1}{\gamma k_{1}}\ln((\bar{\rho}/2\!-\!\epsilon)/k_{2})$, i.e., the convergence to the set $\{w_{i}\in{\mathbb{R}}^{n}\mid\left\|w_{i}^{t}\!-\!x^{\star}\right\|\leq\bar{\rho}\}$ is linear. $\square$
The proof of Theorem 1 is provided in Section IV-C. Theorem 1 provides a semi-global, practical exponential-stability result restricted to the set $\mathcal{D}_{r}:=\mathcal{D}_{1,r}\times\dots\times\mathcal{D}_{N,r}$. Indeed, it is semi-global because the parameters $\gamma^{\star}$ and $\delta^{\star}$ depend on the initial radius $r$ and it is practical because they also depend on the arbitrary small final radius $\bar{\rho}>0$.
IV Extremum Seeking Tracking:
Stability Analysis
In this section, we analyze Extremum Seeking Tracking. Assumptions 12, and 3 hold throughout the whole section. First, let the coordinates $x_{i}^{t},z_{i}^{t}\in{\mathbb{R}}^{n}$ be defined as
$x_{i}^{t}:=w_{i}^{t}-\delta d_{i}^{t},\quad z_{i}^{t}:=s_{i}^{t}-\frac{2f_{i}(w_{i}^{t})d_{i}^{t}}{\delta},$ (6)
for all $i\in\{1,\ldots,N\}$, which allow us to rewrite (4) as
$x_{i}^{t+1}$ $=x_{i}^{t}-\gamma\bigg{(}\sum_{j\in\mathcal{N}_{i}}{\ell}_{ij}x_{j}^{t}+z_{i}^{t}+\frac{2f_{i}(x_{i}^{t}+\delta d_{i}^{t})d_{i}^{t}}{\delta}\bigg{)}$ (7a)
$z_{i}^{t+1}$ $=z_{i}^{t}-\gamma\sum_{j\in\mathcal{N}_{i}}\ell_{ij}\left(z_{j}^{t}+\frac{2f_{j}(x_{j}^{t}+\delta d_{j}^{t})d_{j}^{t}}{\delta}\right).$ (7b)
Remark 2
The new variables $x_{i}^{t}$ and $z_{i}^{t}$ allow us to interpret Extremum Seeking Tracking as an approximated discrete-time version of the continuous-time gradient tracking method proposed in [38]. Indeed, if the case gradients were available, the distributed algorithm (7) would become
$x_{i}^{t+1}$ $={x_{i}^{t}}-{\gamma}\sum_{j\in\mathcal{N}_{i}}{\ell}_{ij}x_{j}^{t}-\gamma\left(z_{i}^{t}+\nabla f_{i}(x_{i}^{t})\right)$ (8a)
$z_{i}^{t+1}$ $={z_{i}^{t}}-{\gamma}\sum_{j\in\mathcal{N}_{i}}{\ell}_{ij}\left(z_{j}^{t}+\nabla f_{j}(x_{j}^{t})\right).$ (8b)
The analysis required to prove Theorem 1 will also require the investigation of the convergence properties of (8) (see Lemma 3 and Remark 3). $\square$
Then, we aggregate the local updates in (7) obtaining the compact algorithm description
${x}^{t+1}$ $=\,x^{t}-\gamma\left(L{x}^{t}+z^{t}+f_{d}(t,x^{t})\right)$ (9a)
${z}^{t+1}$ $=\,z^{t}-\gamma\left(L{z}^{t}+Lf_{d}(t,x^{t})\right),$ (9b)
where we introduced $L:=\mathcal{L}\otimes I_{n}$, $x^{t}:=\mathtt{col}(x_{1}^{t},\dots,x_{N}^{t})$, $z^{t}:=\mathtt{col}(z_{1}^{t},\dots,z_{N}^{t})$, $d^{t}:=\mathtt{col}(d_{1}^{t},\dots,d_{N}^{t})$, and the function $f_{d}:\mathbb{N}\times{\mathbb{R}}^{Nn}\to{\mathbb{R}}^{Nn}$ defined as
$f_{d}(t,x):=\begin{bmatrix}2f_{1}(x_{1}+\delta d_{1}^{t})d_{1}^{t}/\delta\\ \vdots\\ 2f_{N}(x_{N}+\delta d_{N}^{t})d_{N}^{t}/\delta\end{bmatrix},$ (10)
where we decomposed $x$ according to $x:=\mathtt{col}(x_{1},\dots,x_{N})$ with $x_{i}\in{\mathbb{R}}^{n}$ for all $i\in\{1,\ldots,N\}$. We point out that, see also Fig. 1, system (9) can be conceived as an extremum-seeking scheme with output map $f(x+\delta d^{t})$.
Figure 1: Block scheme of the proposed Extremum Seeking Tracking algorithm in the $(x,z)$ coordinates.
We now give an overview of the main steps of the stability analysis carried out to prove Theorem 1:
  1. (i)
    We perform a first change of variables to describe the dynamics (9) in terms of the mean value (over the agents) $\tilde{z}_{\text{avg}}$ of $z$ and the orthogonal part $\tilde{z}_{\perp}$. Then, by relying on averaging theory (see [39] for discrete-time systems or [40, Ch.10] and [41] for continuous-time ones), we introduce a suitable auxiliary system named averaged system. The latter is obtained by averaging the original algorithm dynamics over a common period. The averaged system is shown to be driven by the cost gradients with additive estimation errors.
  2. (ii)
    When neglecting these errors, the averaged system corresponds to an equivalent form of (8). Based on this observation, we rely on existing stability properties of the continuous gradient tracking to demonstrate that the trajectories of the averaged system exponentially converge to an arbitrarily small neighborhood of $\mathtt{col}(\mathbf{1}x^{\star},\tilde{z}^{\text{eq}}_{\perp})$ for some $\tilde{z}^{\text{eq}}_{\perp}$ arising from the analysis.
  3. (iii)
    Finally, we prove Theorem 1 by exploiting the steps above and by using averaging theory to show the closeness between the trajectories of (9) and those of its averaged system.
Step (i) is performed in Section IV-A, step (ii) is carried out in Section IV-B, while Section IV-C is devoted to step (iii).
IV-A Coordinate changes and averaged system
We start by introducing a change of coordinates to highlight the error dynamics of (9) with respect to $\mathtt{col}(x^{\star},-G(\mathbf{1}x^{\star}))$, i.e., the point in which each $x_{i}^{t}$ coincides with the optimal problem solution $x^{\star}$ and perfect tracking is achieved via $z_{i}^{t}$ (see [38]). To this end, let $G:{\mathbb{R}}^{Nn}\to{\mathbb{R}}^{Nn}$ be defined as
$G(x):=$ $\,\mathtt{col}(\nabla f_{1}(x_{1}),\dots,\nabla f_{N}(x_{N})).$ (11)
Then, let the error coordinates $\tilde{x},\tilde{z}\in{\mathbb{R}}^{Nn}$ be defined as
$\tilde{x}:=x-\mathbf{1}x^{\star},\quad\tilde{z}=z+G(\mathbf{1}x^{\star}),$ (12)
and let us introduce $\phi_{xz}:\mathbb{N}\times{\mathbb{R}}^{Nn}\times{\mathbb{R}}^{Nn}\to{\mathbb{R}}^{2Nn}$ as
$\phi_{xz}(t,\tilde{x},\tilde{z})\!\!:=\!\!\begin{bmatrix}-L\tilde{x}-\tilde{z}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -L\tilde{z}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!,$
Then, by using the new coordinates, we rewrite (9) as
$\begin{bmatrix}\tilde{x}^{t+1}\\ \tilde{z}^{t+1}\end{bmatrix}=\begin{bmatrix}\tilde{x}^{t}\\ \tilde{z}^{t}\end{bmatrix}+\gamma\phi_{xz}(t,\tilde{x}^{t},\tilde{z}^{t}),$ (13)
where we have used the property $L\mathbf{1}=0$. As in [38], we take advantage of the initialization $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$. To this end, we introduce the novel coordinates $z_{\text{m}}\in{\mathbb{R}}^{n}$ and $\tilde{z}_{\perp}\in{\mathbb{R}}^{(N-1)n}$ representing the average of the variables $z_{i}^{t}$ (over the agents) and the orthogonal counterpart, namely
$\begin{bmatrix}\tilde{z}_{\text{m}}\\ \tilde{z}_{\perp}\end{bmatrix}:=\begin{bmatrix}\frac{\mathbf{1}^{\top}}{N}\\ R^{\top}\end{bmatrix}\tilde{z},$ (14)
where we introduced the matrix $R\in{\mathbb{R}}^{Nn\times(N-1)n}$ such that $R^{\top}\mathbf{1}=0$, $R^{\top}R=I$. Further, given $n_{\xi}:=(2N-1)n$, let us also introduce $\xi\in{\mathbb{R}}^{n_{\xi}}$ defined as
$\xi:=\begin{bmatrix}\tilde{x}^{\top}&\tilde{z}_{\perp}^{\top}\end{bmatrix}^{\top}.$ (15)
Since $\mathbf{1}^{\top}L=0$ in light of Assumption 1 and $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=\sum_{i=1}^{N}\nabla f_{i}(x^{\star})=0$, we use (14) and (15) to rewrite (13) as
$\xi^{t+1}$ $=\xi^{t}+\gamma\phi_{\xi}(t,\mathtt{col}(\tilde{x}^{t},\mathbf{1}\tilde{z}_{\text{m}}^{t}+R\tilde{z}_{\perp}^{t}))$ (16a)
$\tilde{z}_{\text{m}}^{t+1}$ $=\tilde{z}_{\text{m}}^{t},$ (16b)
where we introduced $\phi_{\xi}:\mathbb{N}\times{\mathbb{R}}^{2Nn}\to{\mathbb{R}}^{n_{\xi}}$ defined as
$\phi_{\xi}(t,\mathtt{col}(\tilde{x},\tilde{z})):=\begin{bmatrix}I&0\\ 0&R\end{bmatrix}\phi_{xz}(t,\tilde{x},\tilde{z}).$ (17)
The equation (16b) implies that $\tilde{z}_{\text{m}}^{t}=\tilde{z}_{\text{m}}^{0}$ for all $t\geq 0$. Then, since $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$ and $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=0$, it holds $\tilde{z}_{\text{m}}^{0}=0$ which allows us to ignore (16b) and rewrite (16) according to the equivalent, reduced system
$\xi^{t+1}=\xi^{t}+\gamma\phi(t,\xi^{t}),$ (18)
where $\phi:\mathbb{N}\times{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is introduced to compactly describe system (16a) with $\tilde{z}_{\text{m}}^{t}=0$ for all $t\geq 0$, i.e., $\phi$ is defined as
$\phi(t,\xi)$ $:=\phi_{\xi}(t,\mathtt{col}(\tilde{x},R\tilde{z}_{\perp}))$ (19)
=$\stackrel{{\scriptstyle(a)}}{{=}}\!\begin{bmatrix}I&0\\ 0&R\end{bmatrix}\!\!\begin{bmatrix}-L\tilde{x}-R\tilde{z}_{\perp}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -LR\tilde{z}_{\perp}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!\!,$
where in $(a)$ we used the definition of $\phi_{\xi}$ (cf. (17)).
We resort to the averaging theory [39] to analyze the time-varying system (18). As customary when this tool is employed, we introduce an auxiliary scheme typically named averaged system, obtained by averaging the time-varying vector field $\phi(t,\xi)$ over $\tau_{\text{per}}$ samples by freezing the system state. In detail, the averaged system associated with (18) is
[…]
with $\xi_{\text{avg}}^{0}=\xi_{0}$, where $\phi_{\text{avg}}:{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is defined as
$\phi_{\text{avg}}(\xi):=\frac{1}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}\phi(k,\xi),\quad\text{for any }t\geq 0.$ (21)
Notice that, since $\phi(k,\xi)$ is periodic with period $\tau_{\text{per}}$ on the first argument, the function $\phi_{\text{avg}}(\xi)$ (cf. (21)) is time-invariant.
The next lemma shows that the gradient of each $f_{i}$ can be approximated by averaging each block of $f_{d}$ (cf. (10)) over the period $\tau_{\text{per}}$ on the first argument. For this reason, the next lemma will be useful to explicitly write $\phi_{\text{avg}}$ (cf. (21)) in terms of $G(x)$, (i.e., the stack of the gradients $\nabla f_{i}(x_{i})$, see (11)).
Lemma 1 (Gradient estimation)
For all $i\in\{1,\ldots,N\}$, there exists $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for any given $x_{i}\in{\mathbb{R}}^{n}$ and all $t\geq 0$, it holds
$\dfrac{2}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}f_{i}(x_{i}+\delta d_{i}^{k})d_{i}^{k}=\nabla f_{i}(x_{i})+\delta^{2}\ell_{i}(x_{i}).$ (22)
Moreover, given any compact set $\mathcal{S}_{i}\subset{\mathbb{R}}^{n}$, if $\delta\in(0,1]$, there exists $L_{i,\mathcal{S}_{i}}>0$ such that
$\left\|\ell_{i}(x_{i})\right\|\leq L_{i,\mathcal{S}_{i}},$ (23)
for all $x_{i}\in\mathcal{S}_{i}$ and $i\in\{1,\ldots,N\}$. $\square$
The proof of Lemma 1 is in Appendix -A.
We note that the quantity on the right-hand side of (22) is time-invariant because $d_{i}^{k}$ is periodic, with period $\tau_{\text{per}}$, and the left-hand side of (22) is averaged over the period $\tau_{\text{per}}$. Now, let us introduce the function $\ell:{\mathbb{R}}^{Nn}\to{\mathbb{R}}^{Nn}$ stacking all the approximation errors $\ell_{i}(x_{i})$ used in (22), namely
$\ell(x):=\mathtt{col}\left(\ell_{1}(x_{1}),\dots,\ell_{N}(x_{N})\right).$ (24)
Then, by using (22), the definitions of $f_{d}$ in (10), $G(x)$ as the stack of the gradients $\nabla f_{i}(x_{i})$ in (11), and $\ell(x)$ as the stack of the approximation errors $\ell_{i}(x_{i})$ in (24), it holds
$\frac{1}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}f_{d}(k,x)=G(x)+\ell(x),$ (25)
for all $x\in{\mathbb{R}}^{Nn}$ and $t\geq 0$. Hence, by introducing $\xi_{\text{avg}}:=\mathtt{col}(\tilde{x}_{\text{avg}},\tilde{z}_{\perp,_{\text{avg}}})$ and by combining the definition of $\phi_{\text{avg}}$ in (21), $\phi$ in (19), and (25), we obtain
$\phi_{\text{avg}}(\xi_{\text{avg}})$ $=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$
$\hskip 11.38092pt+\begin{bmatrix}-\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\\ -R^{\top}L\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\end{bmatrix}.$ (26)
We define
$\phi_{GT}(\xi_{\text{avg}})$ $:=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$
$B$ $:=\begin{bmatrix}-I\\ -R^{\top}L\end{bmatrix},\quad u(\xi_{\text{avg}}):=\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})$
to short (26) as
$\phi_{\text{avg}}(\xi_{\text{avg}})=\phi_{GT}(\xi_{\text{avg}})+\delta^{2}Bu(\xi_{\text{avg}}),$
which, in turn, allows us to rewrite (20) as
[…]
IV-B Averaged System Analysis
In this subsection, we analyze the averaged system (27). To this end, we first consider an additional nominal system in which the term $\gamma\delta^{2}Bu(\xi_{\text{avg}}^{t})$ (i.e., the term describing the gradients’ estimation error) is neglected. Then, by using such a nominal system analysis as a building block, we provide the result concerning system (27). Therefore, we start by studying
$\xi_{\text{avg}}^{t+1}=\xi_{\text{avg}}^{t}+\gamma\phi_{GT}(\xi_{\text{avg}}^{t}),$ (28)
which corresponds to system (27) in the case of $\gamma\delta^{2}Bu(\xi_{\text{avg}}^{t})=0$. The next lemma proves the global exponential stability of the origin for (28).
Lemma 2
There exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $a_{1},a_{2},c_{1}>0$, and $\gamma_{0}>0$ such that, for any $\gamma\in(0,\gamma_{0})$, along the trajectories of (28) it holds
$a_{1}I\leq P$ $\leq a_{2}I$ (29a)
${\xi_{\text{avg}}^{t+1}}^{\top}P\xi_{\text{avg}}^{t+1}-{\xi_{\text{avg}}^{t}}^{\top}P\xi_{\text{avg}}^{t}$ $\leq-\gamma c_{1}\left\|\xi_{\text{avg}}^{t}\right\|^{2},$ (29b)
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$. $\square$
The proof of Lemma 2 is in Appendix -B.
Remark 3
Notice that Lemma 2 proves the algorithm (8) linearly converges to the minimizer of (1), since (28) is an equivalent formulation of (8). $\square$
With this result at hand, we analyze the impact of $u(\cdot)$ thus obtaining the stability properties for the averaged system (27).
Lemma 3
Consider the averaged system (27). Then, for any $r_{\xi}>0$ and $\rho\in(0,r_{\xi})$, there exist $c_{3}\in(0,c_{1})$ and $\delta^{\star}_{1}\in(0,1]$ such that, for any $\gamma\in\min\{\gamma_{0},1\}$, $\delta\in(0,\delta^{\star}_{1})$, and $\left\|\xi_{\text{avg}}^{0}\right\|\leq r_{\xi}$, it holds
(i) $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\sqrt{a_{2}/a_{1}}r_{\xi}}$ for all $t\geq 0$,
$\|\xi_{\text{avg}}^{t}\|\leq\sqrt{a_{2}/a_{1}}\exp\left(-t\gamma c_{3}\right)\|\xi_{\text{avg}}^{0}\|,$ (30)
for all $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$. $\square$
The proof of Lemma 3 is in Appendix -C.
Remark 4
The result of Lemma 3 only involves the averaged system (27). In the next section, such a result will be used as a building block to study the original dynamics, i.e., system (18) and, thus, to conclude the proof of Theorem 1. However, it is a per se result that can be used to show robust stability for the distributed algorithm (8). $\square$
IV-C Proof of Theorem 1
Despite averaging tools for discrete-time systems are already present in the literature, see [39] for example, we got the inspiration from continuous-time averaging [40, Ch. 10] and [41] for elaborating the proof of Theorem 1 to make clear how $\gamma$ affects the closeness of the trajectories of (18) and (27). Since Assumptions 12, and 3 hold, we apply Lemma 2 to claim that there exists $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $\gamma_{0}>0$, and $a_{1},a_{2},c_{1}>0$ such that, if $\gamma\in(0,\gamma_{0})$, the conditions (29) are satisfied. Then, we evaluate the norm of the initial conditions of system (18) and (27), i.e., $\left\|\xi^{0}\right\|=\left\|\xi^{0}_{\text{avg}}\right\|$. By using the definition of $\xi$ (cf. (15)), the changes of variables (12) and (14), and the triangle inequality, we get
$\left\|\xi^{0}\right\|$ $\leq\left\|x^{0}-\mathbf{1}x^{\star}\right\|+\left\|R^{\top}(z^{0}+G(\mathbf{1}x^{\star}))\right\|$
$\hskip 11.38092pt+\left\|\frac{\mathbf{1}^{\top}}{N}(z^{0}+G(\mathbf{1}x^{\star}))\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}r\sqrt{N}+\left\|R\right\|\left\|G(\mathbf{1}x^{\star})\right\|,$
where in $(a)$ we combine the initialization $\left\|x_{i}^{0}-x^{\star}\right\|\leq r$ and $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$ with the fact that $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=\sum_{i=1}^{N}f_{i}(x^{\star})=0$. Hence, by defining $r_{\xi}:=r\sqrt{N}+\left\|R\right\|\left\|G(\mathbf{1}x^{\star})\right\|$, we claim that
$\left\|\xi^{0}\right\|=\left\|\xi_{\text{avg}}^{0}\right\|\leq r_{\xi}.$
Once the initial distance from the origin has been evaluated, we choose any $\bar{\rho}>0$, set $c_{2}:=\sqrt{a_{2}/a_{1}}$, and choose any $\epsilon\in(0,\bar{\rho}(1+c_{2})/2)$. Then, we pick $\rho\in(0,(\bar{\rho}/2-(1+c_{2})\epsilon)/c_{2})$, $c_{1}^{\prime}\in(0,c_{1})$, and use the matrix $P$ satisfying (29) to apply Lemma 3. Specifically, we claim that there exist $c_{3}>0$, $\delta^{\star}_{1}\in(0,1]$, and $\gamma_{0}>0$ such that, for any $\delta\in(0,\delta^{\star}_{1})$ and $\gamma\in(0,\min\{\gamma_{0},\}]$, it holds $\xi_{\text{avg}}^{t}\in\mathcal{B}_{c_{2}r_{\xi}}$ for all $t\geq 0$ and
$\left\|\xi_{\text{avg}}^{t}\right\|\leq c_{2}\exp(-t\gamma c_{3})\left\|\xi_{\text{avg}}^{0}\right\|,$ (31)
for all $\xi_{\text{avg}}^{t}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$. Now, we proceed by finding a bound $\gamma_{1}>0$ such that, for any $\gamma\in(0,\gamma_{1})$, we guarantee the $\epsilon$-closeness between the state $\xi_{\text{avg}}^{t}$ of the averaged system (20) and the one of the original system (18), i.e., that $\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon$ holds true for all $t\geq 0$. To this end, let us introduce
$\upsilon(t,\xi_{\text{avg}}):=\sum_{k=0}^{t-1}\left(\phi(k,\xi_{\text{avg}})-\phi_{\text{avg}}(\xi_{\text{avg}})\right).$
By using this definition and the one of $\phi_{\text{avg}}$ (cf. (21)), it holds
$\upsilon(t+1,\xi_{\text{avg}}^{t+1})-\upsilon(t,\xi_{\text{avg}}^{t})$
$=\phi(t,\xi_{\text{avg}}^{t+1})-\phi_{\text{avg}}(\xi_{\text{avg}}^{t+1})+\upsilon(t,\xi_{\text{avg}}^{t+1})-\upsilon(t,\xi_{\text{avg}}^{t}).$ (32)
Then, let $r_{\xi}^{\prime}:=c_{2}r_{\xi}$ and define $\Delta:=\delta\sqrt{Nn}$. Under the assumption of $\xi^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\geq 0$ (later verified by a proper selection of $\gamma$), we claim that the arguments of the functions $f_{i}$ and their derivatives (embedded into the definitions of $\phi(t,\cdot)$ and $\phi_{\text{avg}}(\cdot)$ and their derivatives) lie into the compact set $\mathcal{B}_{r^{\prime}_{\xi}+\epsilon+\Delta}$. Thus, since the functions $f_{i}$ and its derivatives are continuous (cf. Assumption 3) and the functions $\phi(\cdot,\cdot)$ and $\nu(\cdot,\cdot)$ are periodic in the first argument, we define
[…]
$\left\|\frac{\partial\phi_{\text{avg}}(\xi)}{\partial\xi}\right\|,\left\|\frac{\partial\nu(t,\xi)}{\partial\xi}\right\|\bigg{\}}.$ (33)
Consequently, for all $\xi,\xi^{\prime}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ and $t\geq 0$, it holds
$\|\upsilon(t,\xi)\|$ $\leq 2L_{\phi}\tau_{\text{per}}$ (34a)
$\|\phi(t,\xi)-\phi(t,\xi^{\prime})\|$ $\leq L_{\phi}\|\xi-\xi^{\prime}\|$ (34b)
$\|\phi_{\text{avg}}(\xi)-\phi_{\text{avg}}(\xi^{\prime})\|$ $\leq L_{\phi}\|\xi-\xi^{\prime}\|$ (34c)
$\|\upsilon(t,\xi)-\upsilon(k,\xi^{\prime})\|$ $\leq 2L_{\phi}\tau_{\text{per}}\|\xi-\xi^{\prime}\|$ (34d)
$\left\|\phi_{\text{avg}}(\xi)\right\|$ $\leq L_{\phi}.$ (34e)
Let us introduce $\zeta^{t}\in{\mathbb{R}}^{n_{\xi}}$ defined as
$\zeta^{t}:=\xi_{\text{avg}}^{t}+\gamma\upsilon(t,\xi_{\text{avg}}^{t}).$ (35)
Then, it holds
$\xi^{t}-\zeta^{t}$ $=\sum_{k=0}^{t-1}(\xi^{k+1}-\xi^{k})-(\zeta^{k+1}-\zeta^{k}),$
add ±$\pm\gamma\sum_{k=0}^{t-1}(\phi(k,\zeta^{k})+\phi(k,\xi_{\text{avg}}^{k})),$ and use (32) to get
$\xi^{t}-\zeta^{t}=\gamma\sum_{k=0}^{t-1}(\phi(k,\xi^{k})-\phi(k,\zeta^{k}))$
$+\gamma\sum_{k=0}^{t-1}(\phi(k,\zeta^{k})-\phi(k,\xi_{\text{avg}}^{k}))$
$\hskip 11.38092pt-\gamma\sum_{k=0}^{t-1}(\phi(k,\xi_{\text{avg}}^{k+1})-\phi(k,\xi_{\text{avg}}^{k}))$
$+\gamma\sum_{k=0}^{t-1}(\phi_{\text{avg}}(\xi_{\text{avg}}^{k+1})-\phi_{\text{avg}}(\xi_{\text{avg}}^{k}))$
$\hskip 11.38092pt-\gamma\sum_{k=0}^{t-1}(\upsilon(k,\xi_{\text{avg}}^{k+1})-\upsilon(k,\xi_{\text{avg}}^{k})).$
Use (18), (27), and (34) to bound
$\|\xi^{t}-\zeta^{t}\|$ $\leq\gamma L_{\phi}\sum_{k=0}^{t-1}\|\xi^{k}-\zeta^{k}\|+\gamma^{2}L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)t.$ (36)
Apply the discrete Gronwall inequality (see [42, 43]) and
$\sum_{k=0}^{t-1}\gamma L_{\phi}k\exp\left(-\gamma L_{\phi}k\right)\leq\sum_{k=0}^{\infty}\gamma L_{\phi}k\exp\left(-\gamma L_{\phi}k\right)=1$
to further bound (36) as
$\|\xi^{t}-\zeta^{t}\|$ $\leq\gamma^{2}L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)t$
$\hskip 11.38092pt+\gamma L_{\phi}2\left(1+2\tau_{\text{per}}\right)\exp\left(\gamma L_{\phi}t\right).$ (37)
The definition of $\zeta^{t}$ (35) and the triangle inequality lead to
$\|\xi^{t}-\xi_{\text{avg}}^{t}\|$ $\leq\left\|\xi^{t}-\zeta^{t}\right\|+\gamma\left\|v(t,\xi_{\text{avg}}^{t})\right\|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\gamma^{2}L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)t+\gamma 2L_{\phi}\tau_{\text{per}}$
$\hskip 11.38092pt+\gamma L_{\phi}2\left(1+2\tau_{\text{per}}\right)\exp\left(\gamma L_{\phi}t\right).$ (38)
where $(a)$ uses (37) to bound the first term and (34a) to bound the second one. Then, set $\theta^{\star}\in$ such that
$\theta^{\star}\geq-\tfrac{1}{c_{3}}\ln\left(\tfrac{(\bar{\rho}/2-\epsilon)/c_{2}}{c_{2}r_{\xi}}\right).$ (39)
Let $\gamma_{2}:=\tfrac{\epsilon/3}{L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)\theta^{\star}}$, $\gamma_{3}:=\tfrac{\epsilon/3}{2L_{\phi}(1+2\tau_{\text{per}})\exp(L_{\phi}\theta^{\star})}$,$\gamma_{4}:=\tfrac{\epsilon/3}{2L_{\phi}\tau_{\text{per}}}$, $\gamma_{1}:=\min\{\gamma_{0},\gamma_{2},\gamma_{3},\gamma_{4},1\}$. Pick $\gamma\in(0,\gamma_{1})$ such that $t^{\star}:=\frac{\theta^{\star}}{\gamma}\in{\mathbb{N}}$. This can be done without loss of generality since $\theta^{\star}$ is a design parameter. Then, we bound (38) as
$\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon,$ (40)
for all $t\in[0,t^{\star}]$. As a consequence, since $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}}$ for all $t\geq 0$, it holds ${\xi}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in[0,t^{\star}]$, i.e., we have verified that the bounds (34) can be used into the interval $[0,t^{\star}]$. Moreover, the exponential law (31) and the expression of $\theta^{\star}$ (cf. (39)) ensure that it holds
$\|\xi_{\text{avg}}^{t}\|\leq(\bar{\rho}/2-\epsilon)/c_{2},$ (41)
for all $t\geq t^{\star}$. Now, by using the triangle inequality, we write
$\|\xi^{t^{\star}}\|\leq\|\xi^{t^{\star}}-\xi_{\text{avg}}^{t^{\star}}\|+\|\xi_{\text{avg}}^{t^{\star}}\|\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\bar{\rho}}{2c_{2}},$ (42)
where in $(a)$ we combined (40) and (41). The inequality (42) guarantees that $\xi^{t^{\star}}\in\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$, hence we proved that the trajectories of (27) enters into $\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$ with linear rate. Next, in order to show that $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\geq t^{\star}$, we divide the set of natural numbers in intervals as $\mathbb{N}=[0,\,t^{\star}]\cup[t^{\star},2t^{\star}]\cup\dots$. Define $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ as the solution to (27) for $\xi_{\text{avg}}^{0}=\xi^{t^{\star}}$ and $k\in[0,t^{\star}]$. Thus, at the beginning of each interval $[t^{\star},2t^{\star}]$, the initial condition of (27) coincides with the one of $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ and lies into $\mathcal{B}_{\bar{\rho}/2}\subseteq\mathcal{B}_{r_{\xi}}$. Thus, we apply the same arguments above to guarantee that
$\|\xi^{k+t^{\star}}-\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})\|\leq\epsilon,$
for all $\gamma\in(0,\gamma^{\star})$, $k\in[0,t^{\star}]$, and $\psi_{\text{avg}}(2t^{\star},\xi^{t^{\star}})\in\mathcal{B}_{(\bar{\rho}/2-\epsilon)/c_{2}}$. By using Lemma 3, we guarantee that system (27) cannot escape from $\mathcal{B}_{\bar{\rho}/2-\epsilon}$, namely $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\bar{\rho}/2-\epsilon}$ for all $t\geq t^{\star}$. Thus, we get $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\in[t^{\star},2t^{\star}]$. By recursively applying the same arguments above for each interval $[jt^{\star},(j+1)t^{\star}]$ with $j=2,3,\dots$ and using $\left\|x_{i}^{t}-x^{\star}\right\|\leq\left\|\xi^{t}\right\|$ for all $i\in\{1,\ldots,N\}$ and $t\geq 0$, we get
$\left\|x_{i}^{t}-x^{\star}\right\|\leq\bar{\rho}/2,$ (43)
for all $i\in\{1,\ldots,N\}$ and $t\geq 0$. The change of coordinates (6), (43), and the triangle inequality lead to
$\left\|w_{i}^{t}-x^{\star}\right\|$ $\leq\frac{\bar{\rho}}{2}+\delta\left\|d_{i}^{t}\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}\bar{\rho}/2+\delta\sqrt{n},$ (44)
where in $(a)$ we use the boundedness of the dither signals. The proof follows from (44) by setting $\delta^{\star}:=\min\left\{\delta^{\star}_{1},\frac{\bar{\rho}}{2\sqrt{n}}\right\}$.
V Numerical Computations on Distributed personalized optimization
To corroborate the theoretical analysis, in this section, we provide numerical computations for the proposed distributed algorithm on a personalized optimization framework.
In several engineering applications, a problem of interest consists of optimizing a performance metric while keeping into account user discomfort terms [9, 44]. In these scenarios, the user discomfort term is usually not known in advance but can be only accessed by measurements. Specifically, we associate to each agent $i\in\{1,\ldots,N\}$ a cost function in the form $f_{i}(w)=w^{\top}Q_{i}w+r_{i}^{\top}w+\log(\sum_{\ell=1}^{n}a_{i\ell}e^{b_{i\ell}w_{\ell}})$ with $Q_{i}=Q_{i}^{\top}\in^{n\times n}$, $r_{i}\in^{n}$ and $a_{i\ell},b_{i\ell}$, for all $\ell\in\{1,\ldots,n\}$. In the following, we provide different sets of simulations to study in detail the features of Extremum Seeking Tracking. In particular, each set consists of Monte Carlo simulations over $20$ randomly generated scenarios in which each one differs from the other in terms of cost functions, communication graphs, and algorithmic variables’ initialization. In particular, unless differently stated, in each trial the agents communicate according to Erdős-Rényi random graphs (see, e.g., [45]) with edge probabilities equal to $0.2$. As for the problem parameters, for each trial and all $i\in\{1,\ldots,N\}$, we generate each matrix $Q_{i}$ by pre- and post-multiplying a diagonal matrix (whose diagonal elements are randomly extracted from the interval $[10^{-3},5\cdot 10^{-3}]$ with uniform probability) with an orthonormal matrix (randomly generated by extracting its elements from the interval $[0,1]$ with uniform probability) and its transpose, respectively. Further, for all $i\in\{1,\ldots,N\}$ and $\ell\in\{1,\ldots,n\}$, we randomly extract the components of $r_{i}$ within the interval $[-10^{-2},3\cdot 10^{-2}]$ and the parameters $a_{i\ell},b_{i\ell}$ within the interval $[0,10^{-3}]$ with a uniform probability. In each set of simulations, we choose the parameters $\tau_{i_{p}}$ and $\phi_{i_{p}}$ according to the following procedure. For all $i\in\{1,\ldots,N\}$, we take $\phi_{i_{p}}=\dfrac{\pi}{4}\left(1+(-1)^{p}\right)$ for all $p=1,\dots,n$, while $\tau_{i_{p}}$ have been chosen as the first $\lfloor(n+1)/2\rfloor$ elements of odd numbers greater than $3$ since it is possible to show that such a set of frequencies satisfy (3). Roughly speaking, $\phi_{i,p}=0$ and $\phi_{i,p}=\pi/2$ are used to create orthogonal functions with the same frequency. It is important to note that this selection is not unique. As for the algorithm parameters $\gamma$ and $\delta$, due to the complexity of the dependencies of the bounds $\gamma^{\star}$ and $\delta^{\star}$ provided in Theorem 1, we choose them via a trial-and-error procedure. In each set of simulations, the performance is evaluated by providing graphical results involving the relative errors ${|\sum_{i}f_{i}(\bar{x}^{t})-f(x^{\star})|}/{|f(x^{\star})|}$ and ${\|\bar{x}^{t}-x^{\star}\|}/{\|x^{\star}\|}$ achieved along the trials of the Monte Carlo simulations, where $\bar{x}^{t}:=\frac{1}{N}\sum_{i=1}^{N}x_{i}^{t}$. Simulations are performed using DISROPT [46], a Python package based on MPI to encode and simulate distributed optimization algorithms.
V-A Monte Carlo simulations for large-scale problems
First, we perform numerical simulations over large-scale problems with networks made of $N=250$ agents. We consider different optimization variable sizes, namely $n=10,20$. Part of these performances has been run on the Marconi100 HPC Cluster of the Italian Cineca. We used $10$ nodes of the cluster and, for each node, we used $25$ cores and $4$ GPUs. The code has been adapted in order to perform part of the computation directly on GPUs. The results are shown in Fig. 2. In detail, increasing the number of agents increases the Lipschitz constant of the system to be averaged. Moreover, a larger domain of initial conditions also implies a potentially larger $L_{\phi}$ constant (cf. (33)). This implies smaller $\gamma^{\star}$, which, fixed the other parameters, makes the convergence slower. The decision variable dimension instead impacts the selection of the dither signal. A larger number of states implies a larger number of frequencies. This, in turn, means a longer time to estimate the gradient (cf. Lemma 1). Notice that, however, the accuracy of the final estimate is guaranteed by design. Indeed, since $\delta$ and $\gamma$ are designed on $\bar{\rho}$, the trajectories of (5) converge to a ball of radius $\bar{\rho}$ independently of the problem size.
Figure 2: Monte Carlo simulations for large scale problems: mean and $1$-standard deviation band of cost error (left) and variable error (right).
V-B Monte Carlo simulations varying number of agents
Second, we test Extremum Seeking Tracking in a framework with $n=10$ and different number of agents, i.e., $N=5,10,20,30$. For each value of $N$, we generate communication graphs with a diameter $d$ such that the ratio $N/d$ is constant while varying $N$. Results are depicted in Fig. 3 show that the algorithm slows down as the number of agents increases.
Figure 3: Monte Carlo simulations for varying number of agents: mean and $1$-standard deviation band of cost error (left) and variable error (right).
V-C Monte Carlo simulations varying the size $n$
Then, we test the algorithm features with a fixed number of agents $N=10$ and different decision variable dimensions $n=1,10,100$. The achieved results are reported in Fig. 4 and show that Algorithm 1 slows down as $n$ increases.
Figure 4: Monte Carlo simulations varying $n$: mean and $1$-standard deviation band of cost error (left) and variable error (right).
We conclude this part by providing the results of a single trial in the case with $n=1$ to show the evolution of the solution estimate of each agent in error coordinate, with respect to the optimal solution to the problem.
Figure 5: Evolution of the agents’ estimates $w_{i}^{t}$ in coordinate error with respect to the optimal solution $x^{\star}$.
V-D Monte Carlo simulations varying the parameter $\delta$
Third, we provide numerical simulations in which we vary $\delta$ in the case in which $N=n=10$. As one may expect from Theorem 1, we provide Fig. 6 to show that the final accuracy of the algorithm increases with smaller values of $\delta$.
Figure 6: Monte Carlo simulations for varying amplitude $\delta$: mean and $1$-standard deviation band of cost error (left) and variable error (right).
V-E Monte Carlo simulations for comparisons
We now perform simulations with $n=30$ and $N=10$ to compare our method with the 1-Point Distributed Stochastic Gradient-Tracking Method (1P-DSGT) by [19]. We remark that, similarly to our scheme, when running 1P-DSGT each agent estimates the local gradient with one query of the objective function at each iteration. These results are depicted in Fig. 7 and have been obtained by running the considered distributed schemes with the same objective functions, communication graphs, and initial conditions of the solution estimates.
Figure 7: Monte Carlo simulations for comparison between Extremum Seeking Tracking and 1P-DSGT in [19]: mean and $1$-standard deviation band of cost error (left) and variable error (right).
In detail, Fig. 7 shows that, although the algorithm by [19] exhibits a faster convergence in the beginning phase of the simulations, our scheme has a better convergence rate and final accuracy. Indeed, although both Extremum Seeking Tracking and 1P-DSGT require a single function query for each iteration, we recall that 1P-DSGT adopts a stochastic dither whose theoretical gradient estimation properties are guaranteed only asymptotically and in the case of ergodic processes. On the opposite, our periodic dither guarantees that the theoretical gradient estimation is (practically) achieved with $\tau_{\text{per}}$ functions’ samples (in an averaged sense). This makes the convergence rate of our algorithm faster.
V-F Monte Carlo simulations in stochastic scenarios
Finally, we compare the considered distributed algorithms in a stochastic scenario in which each agent receives cost measurements affected by noise, i.e., in the case of $f_{i}(w_{i}^{t})+\eta_{i}^{t}$ in place of $f_{i}(w_{i}^{t})$ in Algorithm 1, where each component $\eta_{i}^{t}\in^{n}$ is randomly generated according to the Gaussian distribution with expected value 00 and standard deviation $0.1$. Fig. 8 compares the behavior of Extremum Seeking Tracking and 1P-DSGT by [19] in the case in which $N=30$ and $n=10$. These plots confirm that Extremum Seeking Tracking exhibits faster convergence and greater final accuracy with respect to the considered scheme also in the considered stochastic setup. To interpret these results, we also remark that Theorem 1 ensures (semi-global, practical) stability properties for Extremum Seeking Tracking. Coherently, Fig. 8 shows the typical behavior exhibited by the trajectories of perturbed systems in the neighborhood of (practically) stable equilibria.
Figure 8: Monte Carlo simulations in stochastic scenarios: mean and $1$-standard deviation band of cost error (left) and variable error (right).
VI Conclusions
In this paper, we addressed a distributed optimization problem in which the cost function is unknown and agents have only access to local measurements. Taking inspiration from a continuous gradient tracking algorithm, we proposed a novel gradient-free distributed optimization algorithm in which gradients are estimated via extremum seeking. We analyzed the convergence properties of the proposed algorithm by using Lyapunov and averaging tools from system theory. We corroborated the theoretical analysis through Monte Carlo simulations on personalized optimization problems.
-A Proof of Lemma 1
Given $\alpha=\mathtt{col}(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}$, $y=\mathtt{col}(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$, and a smooth function $f\,:\,\mathbb{R}^{n}\to\mathbb{R}$, we define
$\alpha!$ $:=\alpha_{1}!\dots\alpha_{n}!,\quad$ $y^{\alpha}:=y_{1}^{\alpha_{1}}\dots y_{n}^{\alpha_{n}},$
$\partial^{\alpha}f(y)$ $:=\dfrac{\partial^{\alpha_{1}}}{\partial y_{1}^{\alpha_{1}}}\dots\dfrac{\partial^{\alpha_{n}}}{\partial y_{n}^{\alpha_{n}}}f(y),\quad$ $|\alpha|:=\alpha_{1}+\dots+\alpha_{n}.$
Being each function $f_{i}$ smooth (cf. Assumption 3), we can apply Taylor’s expansion (cf. [47, Theorem 2]) and write
$f_{i}(x_{i}+\delta d_{i}^{t})\!$ $=\!f_{i}(x_{i})\!+\!\delta{d_{i}^{t}}^{\top}\nabla f_{i}(x_{i})\!+\!\frac{\delta^{2}}{2}{d_{i}^{t}}^{\top}\nabla^{2}f_{i}(x_{i})d_{i}^{t}$
$\hskip 11.38092pt+\delta^{3}R_{i,2}(x_{i},\delta d_{i}^{t}),$ (45)
where the remainder $R_{i,2}(x_{i},\delta d_{i}^{t})$ is given by
$R_{i,2}(x_{i},\delta d_{i}^{t})=\sum_{|\alpha|=3}\dfrac{\partial^{\alpha}f_{i}(x_{i}+c\delta d_{i}^{t})}{\alpha!}(\delta d_{i}^{t})^{\alpha},$ (46)
for some $c\in(0,1)$. Then, we can use (45) to write
$\dfrac{2}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}f_{i}(x_{i}+\delta d_{i}^{k})$
$=\dfrac{2f_{i}(x_{i})}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}+\left[\dfrac{2}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}\left(d_{i}^{k}{d_{i}^{k}}^{\top}\right)\right]\nabla f_{i}(x_{i})$
$\hskip 11.38092pt+\dfrac{\delta}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}\left(d_{i}^{k}{d_{i}^{k}}^{\top}\right)\nabla^{2}f_{i}(x_{i})d_{i}^{k}$
$\hskip 11.38092pt+\dfrac{2}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}R_{i,2}(x_{i},\delta d_{i}^{k}).$ (47)
By combining (3) and (47), we get
$\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}=0,\quad\dfrac{2}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}\left(d_{i}^{k}{d_{i}^{k}}^{\top}\right)=I_{n}$
$\sum_{k=t+1}^{t+\tau_{\text{per}}}\left(d_{i}^{k}{d_{i}^{k}}^{\top}\right)\nabla^{2}f_{i}(x_{i})d_{i}^{k}=0,$
which combined with (47), allow us to write
$f_{i}(x_{i}+\delta d_{i}^{t})=\nabla f_{i}(x_{i})+\dfrac{2}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}R_{i,2}(x_{i},\delta d_{i}^{k}).$
The proof follows by setting $\ell_{i}(x_{i})=\dfrac{2}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}R_{i,2}(x_{i},\delta d_{i}^{k})/\delta^{3}$. Finally, given a compact set $\mathcal{S}_{i}\subset{\mathbb{R}}^{n}$, let us bound $\left\|\ell_{i}(x_{i})\right\|$ for all $x_{i}\in\mathcal{S}_{i}$. Note that $\left\|\delta d_{i}^{t}\right\|\leq\delta\sqrt{n}$ for all $t\geq 0$ and let $\mathcal{S}^{\prime}_{i}\subset{\mathbb{R}}^{n}$ be a compact set such that (i) $\mathcal{S}_{i}\subseteq\mathcal{S}_{i}^{\prime}\subset{\mathbb{R}}^{n}$, and (ii) $x_{i}+\delta d_{i}^{t}\in\mathcal{S}^{\prime}$ for all $x_{i}\in\mathcal{S}_{i}$, $\delta\in(0,1]$, and $t\geq 0$. Thus, we can write
$\sup_{\begin{subarray}{c}x_{i}\in\mathcal{S}_{i}^{\prime}\\ k\in\{1,\ldots,\tau_{\text{per}}-1\}\end{subarray}}\left\|\frac{R_{i,2}(x_{i},\delta d_{i}^{k})}{\delta^{3}}\right\|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sup_{\begin{subarray}{c}x_{i}\in\mathcal{S}_{i}^{\prime}\\ k\in\{1,\ldots,\tau_{\text{per}}-1\}\end{subarray}}\left\|\frac{1}{\delta^{3}}\sum_{|\alpha|=3}\dfrac{\partial^{\alpha}f_{i}(x_{i})}{\alpha!}(\delta d_{i}^{k})^{\alpha}\right\|$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{\begin{subarray}{c}x_{i}\in\mathcal{S}_{i}^{\prime}\\ k\in\{1,\ldots,\tau_{\text{per}}-1\}\end{subarray}}\left\|\sum_{|\alpha|=3}\dfrac{\partial^{\alpha}f_{i}(x_{i})}{\alpha!}(d_{i}^{k})^{\alpha}\right\|=:L_{i,\mathcal{S}_{i}}^{\prime},$ (48)
where in $(a)$ we use the expression (46) of $R_{i,2}(x_{i},\delta d_{i}^{k})$, the definition of $\mathcal{S}_{i}^{\prime}$, and the fact that $\delta\in(0,1]$, while in $(b)$ we drop out the term $\delta^{3}$ from $(\delta d_{i}^{k})^{\alpha}$. We underline that, since the set $\mathcal{S}_{i}^{\prime}$ is compact and $f_{i}$ is smooth, $L_{i,\mathcal{S}_{i}}^{\prime}$ exists and is finite. The bound of $\ell_{i}(x_{i})$ follows by defining $L_{i,\mathcal{S}_{i}}:=2L_{i,\mathcal{S}_{i}}^{\prime}\sqrt{n}$ and combining the result (48) with the bound about the norm of the dither signal, i.e., $\left\|d_{i}^{t}\right\|\leq\sqrt{n}$ for all $t\geq 0$.
-B Proof of Lemma 2
In [38], it is provided a Lyapunov function proving that, under the Assumptions 12, and 3, the point $\xi^{\star}=(\mathbf{1}x^{\star},\tilde{z}_{\perp}^{\text{eq}})$, with $\tilde{z}_{\perp}^{\text{eq}}:=-R^{\top}G(\mathbf{1}x^{\star})$, is a globally exponentially stable equilibrium for the continuous-time system
$\dot{\xi}(t)=\phi_{\text{GT}}(\xi(t)).$
In detail, [38, Th. 3.1] introduces $\bar{T}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ to define $\bar{\xi}:=\bar{T}\xi$, and a matrix $\bar{P}=\bar{P}^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ such that
$\bar{a}_{1}I$ $\leq\bar{P}\leq\bar{a}_{2}I$ (49a)
$2\bar{\xi}_{\text{avg}}^{\top}\bar{P}\bar{T}\phi_{\text{GT}}(\bar{T}^{-1}\bar{\xi}_{\text{avg}})$ $\leq-\bar{a}_{3}\left\|\bar{\xi}_{\text{avg}}\right\|^{2},$ (49b)
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$ and some $\bar{a}_{1},\bar{a}_{2},\bar{a}_{3}>0$. Let $P:=\bar{T}^{-1}\bar{P}T$. Then, there exist $a_{1},a_{2},a_{3}>0$ such that
$a_{1}I$ $\leq P\leq a_{2}I$ (50a)
$2\xi_{\text{avg}}^{\top}P\phi_{\text{GT}}(\xi_{\text{avg}})$ $\leq-a_{3}\left\|\xi_{\text{avg}}\right\|^{2},$ (50b)
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$. Hence, we define the candidate Lyapunov function $V(\xi_{\text{avg}}):=\xi_{\text{avg}}^{\top}P\xi_{\text{avg}}$ and bound $\Delta V(\xi_{\text{avg}}^{t}):=V(\xi_{\text{avg}}^{t+1})-V(\xi_{\text{avg}}^{t})$ along the trajectories of (28) as
$\Delta V(\xi_{\text{avg}}^{t})\leq-\gamma a_{3}\left\|\xi_{\text{avg}}^{t}\right\|^{2}+\gamma^{2}\phi_{\text{GT}}(\xi_{\text{avg}}^{t})^{\top}P\phi_{\text{GT}}(\xi_{\text{avg}}^{t}).$ (51)
Moreover, by using the Lipschitz continuity of the gradients of the objective functions (cf. Assumption 3) and the definition of $\phi_{\text{GT}}$, there exists $a_{4}>0$ such that
$\left\|\phi_{\text{GT}}(\xi_{\text{avg}}^{t})\right\|\leq a_{4}\left\|\xi_{\text{avg}}^{t}\right\|.$ (52)
Finally, for any $c_{1}\in(0,a_{3})$, let $\gamma_{0}:=(a_{3}-c_{1})/(a_{2}a_{4}^{2})$ and the proof follows by using (51) and (52).
-C Proof of Lemma 3
The proof relies on (i) the matrix $P$ satisfying (29), and (ii) the fact that the norm of the perturbation term $\gamma\delta Bu(\xi_{\text{avg}}^{t})$ can be arbitrarily reduced through the parameter $\delta$ as long as $\xi_{\text{avg}}^{t}$ lies into a compact set. First of all, without loss of generality, we assume $\rho\leq r_{\xi}$. Indeed, we will use the parameter $r_{\xi}$ to define a (compact) ball and arbitrarily bound the norm of the perturbation term $\gamma\delta Bu(\xi_{\text{avg}}^{t})$ through the parameters $\delta$ as long as $\xi_{\text{avg}}^{t}$ lies into this ball. Hence, we can always use the more conservative condition. In detail, we introduce the candidate Lypaunov function $V(\xi_{\text{avg}}):=\xi_{\text{avg}}^{\top}P\xi_{\text{avg}}$ and the set $\Omega_{r_{\xi}}:=\{\xi_{\text{avg}}\in^{n_{\xi}}\mid V(\xi_{\text{avg}})\leq a_{2}r_{\xi}^{2}\}\subset{\mathbb{R}}^{n_{\xi}}$. Then, from (29a), we derive $\mathcal{B}_{r_{\xi}}\subseteq\Omega_{r_{\xi}}\subseteq\mathcal{B}_{r^{\prime}_{\xi_{\text{avg}}}}$, where $r^{\prime}_{\xi_{\text{avg}}}:=\sqrt{a_{2}/a_{1}}r_{\xi}$. Now, under the assumption $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}}$ (later verified by a proper selection of the algorithm parameters), it holds $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$. By using this property and since $\gamma\in(0,\gamma_{0})$, we use (29b), the Cauchy-Schwarz inequality, and (52) to bound $\Delta V(\xi_{\text{avg}}^{t}):=V(\xi_{\text{avg}}^{t+1})-V(\xi_{\text{avg}}^{t})$ along the trajectories of (27) as
$\Delta V(\xi_{\text{avg}}^{t})$ $\leq-\gamma c_{1}\|\xi_{\text{avg}}^{t}\|^{2}+\gamma\delta^{2}2\left\|PB\right\|\|\xi_{\text{avg}}^{t}\|\left\|u(\xi_{\text{avg}}^{t})\right\|$
$\hskip 11.38092pt+\delta^{2}\gamma^{2}2a_{4}\left\|PB\right\|\left\|\xi_{\text{avg}}^{t}\right\|\left\|u(\xi_{\text{avg}}^{t})\right\|$
$\hskip 11.38092pt+\delta^{4}\gamma^{2}\left\|B^{\top}PB\right\|\left\|u(\xi_{\text{avg}}^{t})\right\|^{2}.$ (53)
Now, for all $i\in\{1,\ldots,N\}$, let us introduce the $\mathcal{S}_{i}:=\{x_{i}\in{\mathbb{R}}^{n}\mid\left\|x_{i}-x^{\star}\right\|\leq r^{\prime}_{\xi_{\text{avg}}}\}$. Hence, we note that $\xi_{\text{avg}}:=(\tilde{x}_{\text{avg}},\tilde{z}_{\perp,\text{avg}})\in\Omega_{r_{\xi}}\implies\tilde{x}\in\mathcal{S}\subset{\mathbb{R}}^{Nn}$, where $\mathcal{S}:=\mathcal{S}_{1}\times\dots\times\mathcal{S}_{N}$. Then, we apply result (23) to claim that, for all $i\in\{1,\ldots,N\}$, it holds $\ell_{i}(x_{i})\leq L_{i,\mathcal{S}_{i}}$ for all $x_{i}\in\mathcal{S}_{i}$. Thus, by defining $L_{\mathcal{S}}:=\max_{i}\{L_{1,\mathcal{S}_{1}},\dots,L_{N,\mathcal{S}_{N}}\}$ and using the definition $u(\xi_{\text{avg}}^{t})=\ell(\tilde{x}^{t}+\mathbf{1}x^{\star})$, we get
$\left\|u(\xi_{\text{avg}}^{t})\right\|\leq\sqrt{N}L_{\mathcal{S}}.$ (54)
Since $\gamma\in(0,1]$ and $\delta\in(0,1]$, we bound (53) as
$\Delta V(\xi_{\text{avg}}^{t})$ $\leq-\gamma c_{1}\|\xi_{\text{avg}}^{t}\|^{2}+\gamma\delta^{2}(b_{1}\left\|\xi_{\text{avg}}^{t}\right\|+b_{2}),$ (55)
where we introduced
$b_{1}$ $:=2\left\|PB\right\|\sqrt{N}L_{\Omega_{r_{\xi}}}+2a_{4}\left\|PB\right\|\sqrt{N}L_{\mathcal{S}}$
$b_{2}$ $:=\left\|B^{\top}PB\right\|NL_{\mathcal{S}}^{2}.$
Therefore, for any $\rho\in(0,r_{\xi})$ and $c_{1}^{\prime}\in(0,c_{1})$, we define
$\delta^{\star}_{1}:=\min\left\{\sqrt{(c_{1}-c_{1}^{\prime})\rho^{2}/(b_{1}r^{\prime}_{\xi_{\text{avg}}}+b_{2})},1\right\}.$ (56)
Hence, by combining (55) and (56), if $\delta\in(0,\delta^{\star}_{1})$, then, for all $\xi_{\text{avg}}^{t}\in\Omega_{r^{\prime}_{\xi_{\text{avg}}}}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$, it holds
$\Delta V(\xi_{\text{avg}}^{t})<-\gamma c_{1}^{\prime}\left\|\xi_{\text{avg}}^{t}\right\|^{2}.$ (57)
Thus, the inequality (57) ensures that the set $\Omega_{r_{\xi}}$ is invariant for system (27). Hence, if we pick $\xi_{\text{avg}}^{0}\in\mathcal{B}_{r_{\xi}}$, we prove that $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$ for all $t\geq 0$. Consequently, the bound (54) holds for all $t\geq 0$ and, in turn, also the inequality (57) is verified for all $t\geq 0$, namely we proved that the trajectories of system (27) enter the ball $\mathcal{B}_{\rho}$ exponentially fast. The result (30) follows from the inequality (57) and (29a) by setting $c_{3}:=c_{1}^{\prime}/(2a_{2})$.

### v6

@@
 We consider the challenging scenario in which objective functions are unknown and agents have only access to local measurements of their local functions.
-We propose a novel distributed algorithm that combines a recent gradient tracking policy with an extremum-seeking technique to estimate the global descent direction.
+We propose a novel distributed algorithm that combines a recent gradient tracking policy with an extremum seeking technique to estimate the global descent direction.
 The joint use of these two techniques results in a distributed optimization scheme that provides arbitrarily accurate solution estimates through the combination of Lyapunov and averaging analysis approaches with consensus theory.
@@
 We organize the literature review in two blocks: collaborative/distributed extremum seeking and distributed zeroth-order/derivative-free optimization.
-Early references on distributed extremum for the so-called consensus optimization framework are [8] for a discrete-time setting and [9] for a continuous-time one.
-A proportional-integral extremum-seeking design technique is proposed in [10], while, in [11], the gradient is approximated through a real-time protocol.
+Early references on distributed extremum seeking for the so-called consensus optimization framework are [8] for a discrete-time setting and [9] for a continuous-time one.
+A proportional-integral extremum seeking design technique is proposed in [10], while, in [11], the gradient is approximated through a real-time protocol.
 In [12], authors propose the use of the sliding mode to generate the dither signal at the base of the extremum seeking.
@@
 In [16], instead, a distributed, continuous-time scheme based on sign-based consensus is designed.
-In [17], a Lie bracket technique and extremum-seeking are used for problems with linear constraints.
-In [18], resource allocation problems are addressed by an extremum-seeking scheme.
-As for distributed zeroth-order/derivative-free optimization, authors in [19] develop a zeroth-order scheme based on a one-point estimator and a gradient tracking policy.
+In [17], a Lie bracket technique and extremum seeking are used for problems with linear constraints.
+In [18], resource allocation problems are addressed by an extremum seeking scheme.
+As for distributed zeroth-order/derivative-free optimization, authors in [19] develop a zeroth-order scheme based on a 1-point estimator and a gradient tracking policy.
 The work [20] instead proposes a zeroth-order algorithm based on a two-point estimator with a distributed gradient descent strategy and another one based on an $n$-point estimator with a gradient tracking policy.
@@
 All the cited papers but [30] prove stability of the proposed schemes by using arguments not based on Lyapunov theory.
-We also note that [30] proposes an algorithm neither based on gradient tracking nor using extremum-seeking.
+We also note that [30] proposes an algorithm neither based on gradient tracking nor using extremum seeking.
 This paper proposes a novel algorithm to solve a distributed optimization problem in which network agents can only evaluate their local cost function at a given point, but not its gradient.
@@
 The convergence of our scheme is proved through Lyapunov and averaging theories for discrete-time systems.
-It is worth mentioning that our scheme, together with the ones in [34, 35], is the only distributed extremum-seeking scheme proposed in discrete-time with the following distinctive features.
+It is worth mentioning that our scheme, together with the ones in [34, 35], is the only distributed extremum seeking scheme proposed in discrete-time with the following distinctive features.
 The work [34] (i) does not address a consensus optimization problem and (ii) relies on consensus dynamics estimating the global cost, while ours estimates the global gradient.
-Instead, in [35], (i) the addressed consensus optimization problems have scalar decision variables and (ii) an extremum-seeking technique is combined with a distributed gradient algorithm, i.e., without a tracking mechanism.
-As for the literature on distributed zeroth-order/derivative-free methods, the closest work is [19] in which extremum seeking is not used and the global gradient is approximated via a randomized one-point policy.
-Consistent with the comparison table provided in [19], we highlight that there are no other distributed algorithms in the literature using 1-point gradient estimators.
+Instead, in [35], (i) the addressed consensus optimization problems have scalar decision variables and (ii) an extremum seeking technique is combined with a distributed gradient algorithm, i.e., without a tracking mechanism.
+As for the literature on distributed zeroth-order/derivative-free methods, the closest work is [19] in which extremum seeking is not used and the global gradient is approximated via a randomized 1-point policy.
+Consistently with the comparison table provided in [19], we highlight that there are no other distributed algorithms in the literature using 1-point gradient estimators.
 Indeed, although the distributed scheme in [36] uses 1-point gradient estimators, we remark that it is tailored for partition-based optimization, i.e., a simplified setup in which each local function depends only on the neighbors’ decision variables, see [3].
@@
 Assumption 3
-Each cost $f_{i}$ is $\mathcal{C}^{3}$ and has $\overline{L}_{i}$-Lipschitz continuous gradients. We denote $\overline{L}=\max\{\overline{L}_{1},\dots,\overline{L}_{N}\}$. $\square$
+Each cost $f_{i}$ is $\mathcal{C}^{3}$ and has $\overline{L}_{i}$-Lipschitz continuous gradients. We denote $\overline{L}:=\max\{\overline{L}_{1},\dots,\overline{L}_{N}\}$. $\square$
 Assumption 2 ensures the existence of a unique solution $x^{\star}\in\mathbb{R}^{n}$ to problem (1).
@@
 Namely, given the iteration index $t\in\mathbb{N}$ and by denoting with $w_{i}^{t}\in{\mathbb{R}}^{n}$ the $i$-th agent’s estimate, at iteration $t$, of the solution to problem (1), our goal is to design a distributed protocol able to steer all these estimates to the minimizer $x^{\star}$.
-The peculiar challenge of this paper is that each agent $i$ can access neither the gradients $\nabla f_{i}(w_{i}^{t})$ (as in standard gradient-based methods) nor the cost functions in arbitrary points (as in standard zeroth-order methods).
+The peculiar challenge of this paper is that each agent $i$ cannot access either the gradients $\nabla f_{i}(w_{i}^{t})$ (as in standard gradient-based methods) or the cost functions in arbitrary points (as in standard zeroth-order methods).
 More in detail, we assume that agent $i$ can only use the single measurement $f_{i}(w_{i}^{t})$ properly combined with the so-called dither signal $d_{i}^{t}\in{\mathbb{R}}^{n}$ and the amplitude parameter $\delta>0$.
@@
 $\sum_{t=0}^{\tau_{\text{per}}-1}\sin\left(\tfrac{2\pi t}{\tau_{i_{p}}}+\phi_{i_{p}}\right)\sin\left(\tfrac{2\pi t}{\tau_{i_{q}}}+\phi_{i_{q}}\right)\!=\!\dfrac{\tau_{\text{per}}}{2}$ (3b)
-$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{r}}}\!+\!\phi_{i_{r}}\right)\!=\!0,$ (3c)
-for $i\in\{1,\ldots,N\}$.
-Here, $\tau_{\text{per}}\in\mathbb{N}$ is the least common multiple of all periods $\tau_{ip}$.
+$\sum_{t=0}^{\tau_{\text{per}}-1}\sin\!\left(\tfrac{2\pi t}{\tau_{i_{p}}}\!+\!\phi_{i_{p}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{q}}}\!+\!\phi_{i_{q}}\right)\!\sin\!\left(\tfrac{2\pi t}{\tau_{i_{r}}}\!+\!\phi_{i_{r}}\right)\!=\!0,$ (3c)
+for all $i\in\{1,\ldots,N\}$.
+Here, $\tau_{\text{per}}\in\mathbb{N}$ is the least common multiple of all periods $\tau_{i_{p}}$.
 Sinusoidal dither functions are useful in practical applications to guarantee smooth inputs to the plant.
-However, other possibilities as, e.g., as square or triangular waves, are possibile [37].
+However, other possibilities, e.g., as square or triangular waves, are possibile [37].
 initialization: $x_{i}^{0}\in\mathbb{R}^{n}$ and $z_{i}^{0}=0$
@@
 First, errors in our method are given by third-order residuals as opposed to second-order ones in finite-difference methods.
-Second, gradient estimation is based on a single function query.
-This could be advantageous in scenarios in which multiple queries are expensive or even not allowed.
+Second, gradient estimation is based on a single-function query per agent.
+This could be advantageous in scenarios in which multiple queries per agent are expensive or even not allowed.
 Third and final, the estimation policy updates, and so the convergence guarantees, are purely deterministic.
@@
 To this end, for all $i\in\{1,\ldots,N\}$ and given $r>0$, we define the set $\mathcal{D}_{i,r}\subset{\mathbb{R}}^{2n}$ as
-$\mathcal{D}_{i,r}\!:=\!\left\{\mathtt{col}(w_{i},z_{i})\in{\mathbb{R}}^{2n}\!\mid\!\left\|w_{i}-x^{\star}\right\|\leq r,s_{i}=2f_{i}(w_{i})/\delta\right\}\!.$
+$\mathcal{D}_{i,r}\!:=\!\left\{\mathtt{col}(w_{i},s_{i})\in{\mathbb{R}}^{2n}\!\mid\!\left\|w_{i}-x^{\star}\right\|\leq r,s_{i}=2f_{i}(w_{i})/\delta\right\}\!.$
 Theorem 1
 Consider (4) and let Assumptions 12, and 3 hold.
-Then, for any $r,\bar{\rho}>0$, there exist $\gamma^{\star},\delta^{\star},k_{1}>0$, $\epsilon\in(0,\bar{\rho}/2)$, and $k_{2}\geq(\bar{\rho}/2-\epsilon)$ such that, for any $\gamma\in(0,\gamma^{\star})$ and $\delta\in(0,\delta^{\star})$, if $\mathtt{col}(w_{i}^{0},s_{i}^{0})\in\mathcal{D}_{i,r}$ for all $i\in\{1,\ldots,N\}$, the trajectories of (4) are bounded and for all $i\in\{1,\ldots,N\}$
+Then, for any $r,\bar{\rho}>0$, there exist $\gamma^{\star},\delta^{\star},k_{1}>0$, $\epsilon\in(0,\bar{\rho}/2)$, and $k_{2}\geq(\bar{\rho}/2-\epsilon)$ such that, for any $\gamma\in(0,\gamma^{\star})$, $\delta\in(0,\delta^{\star})$, and $\mathtt{col}(w_{i}^{0},s_{i}^{0})\in\mathcal{D}_{i,r}$ for all $i\in\{1,\ldots,N\}$, the trajectories of (4) are bounded and satisfy
 $\left\|w_{i}^{t}-x^{\star}\right\|\leq\bar{\rho},$ (5)
-for all $t\geq t^{\star}$, where $t^{\star}\!:=\!-\frac{1}{\gamma k_{1}}\ln((\bar{\rho}/2\!-\!\epsilon)/k_{2})$, i.e., the convergence to the set $\{w_{i}\in{\mathbb{R}}^{n}\mid\left\|w_{i}^{t}\!-\!x^{\star}\right\|\leq\bar{\rho}\}$ is linear.
+for all $i\in\{1,\ldots,N\}$ and $t\geq t^{\star}:=\!-\frac{1}{\gamma k_{1}}\ln((\bar{\rho}/2\!-\!\epsilon)/k_{2})$, i.e., the convergence to the set $\{w_{i}\in{\mathbb{R}}^{n}\mid\left\|w_{i}^{t}\!-\!x^{\star}\right\|\leq\bar{\rho}\}$ is linear.
 $\square$
@@
 where we decomposed $x$ according to $x:=\mathtt{col}(x_{1},\dots,x_{N})$ with $x_{i}\in{\mathbb{R}}^{n}$ for all $i\in\{1,\ldots,N\}$.
-We point out that, see also Fig.
-1, system (9) can be conceived as an extremum-seeking scheme with output map $f(x+\delta d^{t})$.
+We point out that system (9) can be conceived as an extremum seeking scheme with output map $f(x+\delta d^{t})$, see also Fig.
+1.
 Figure 1: Block scheme of the proposed Extremum Seeking Tracking algorithm in the $(x,z)$ coordinates.
@@
   1. (i)
-    We perform a first change of variables to describe the dynamics (9) in terms of the mean value (over the agents) $\tilde{z}_{\text{avg}}$ of $z$ and the orthogonal part $\tilde{z}_{\perp}$.
+    We perform two change of variables to describe the dynamics (9) in terms of the mean value (over the agents) $\tilde{z}_{\text{avg}}$ of $z$ and the orthogonal part $\tilde{z}_{\perp}$ (both in error coordinates).
 Then, by relying on averaging theory (see [39] for discrete-time systems or [40, Ch.10] and [41] for continuous-time ones), we introduce a suitable auxiliary system named averaged system.
@@
 and let us introduce $\phi_{xz}:\mathbb{N}\times{\mathbb{R}}^{Nn}\times{\mathbb{R}}^{Nn}\to{\mathbb{R}}^{2Nn}$ as
-$\phi_{xz}(t,\tilde{x},\tilde{z})\!\!:=\!\!\begin{bmatrix}-L\tilde{x}-\tilde{z}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -L\tilde{z}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!,$
+$\phi_{xz}(t,\tilde{x},\tilde{z})\!\!:=\!\!\begin{bmatrix}-L\tilde{x}-\tilde{z}-f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -L\tilde{z}-L(f_{d}(t,\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}\!.$
 Then, by using the new coordinates, we rewrite (9) as
@@
 As in [38], we take advantage of the initialization $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$.
-To this end, we introduce the novel coordinates $z_{\text{m}}\in{\mathbb{R}}^{n}$ and $\tilde{z}_{\perp}\in{\mathbb{R}}^{(N-1)n}$ representing the average of the variables $z_{i}^{t}$ (over the agents) and the orthogonal counterpart, namely
+To this end, we introduce the novel coordinates $\tilde{z}_{\text{m}}\in{\mathbb{R}}^{n}$ and $\tilde{z}_{\perp}\in{\mathbb{R}}^{(N-1)n}$ representing the average of the variables $\tilde{z}_{i}^{t}$ (over the agents) and the orthogonal counterpart, namely
 $\begin{bmatrix}\tilde{z}_{\text{m}}\\ \tilde{z}_{\perp}\end{bmatrix}:=\begin{bmatrix}\frac{\mathbf{1}^{\top}}{N}\\ R^{\top}\end{bmatrix}\tilde{z},$ (14)
@@
 $\phi_{\xi}(t,\mathtt{col}(\tilde{x},\tilde{z})):=\begin{bmatrix}I&0\\ 0&R\end{bmatrix}\phi_{xz}(t,\tilde{x},\tilde{z}).$ (17)
-The equation (16b) implies that $\tilde{z}_{\text{m}}^{t}=\tilde{z}_{\text{m}}^{0}$ for all $t\geq 0$.
+The equation (16b) implies that $\tilde{z}_{\text{m}}^{t}=\tilde{z}_{\text{m}}^{0}$ for all $t\in\mathbb{N}$.
 Then, since $z_{i}^{0}=0$ for all $i\in\{1,\ldots,N\}$ and $\mathbf{1}^{\top}G(\mathbf{1}x^{\star})=0$, it holds $\tilde{z}_{\text{m}}^{0}=0$ which allows us to ignore (16b) and rewrite (16) according to the equivalent, reduced system
 $\xi^{t+1}=\xi^{t}+\gamma\phi(t,\xi^{t}),$ (18)
-where $\phi:\mathbb{N}\times{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is introduced to compactly describe system (16a) with $\tilde{z}_{\text{m}}^{t}=0$ for all $t\geq 0$, i.e., $\phi$ is defined as
+where $\phi:\mathbb{N}\times{\mathbb{R}}^{n_{\xi}}\to{\mathbb{R}}^{n_{\xi}}$ is introduced to compactly describe system (16a) with $\tilde{z}_{\text{m}}^{t}=0$ for all $t\in\mathbb{N}$, i.e., $\phi$ is defined as
 $\phi(t,\xi)$ $:=\phi_{\xi}(t,\mathtt{col}(\tilde{x},R\tilde{z}_{\perp}))$ (19)
@@
 Lemma 1 (Gradient estimation)
-For all $i\in\{1,\ldots,N\}$, there exists $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for any given $x_{i}\in{\mathbb{R}}^{n}$ and all $t\geq 0$, it holds
+For all $i\in\{1,\ldots,N\}$, there exists $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for any given $x_{i}\in{\mathbb{R}}^{n}$ and all $t\in\mathbb{N}$, it holds
 $\dfrac{2}{\delta\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}f_{i}(x_{i}+\delta d_{i}^{k})d_{i}^{k}=\nabla f_{i}(x_{i})+\delta^{2}\ell_{i}(x_{i}).$ (22)
@@
 $\frac{1}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}f_{d}(k,x)=G(x)+\ell(x),$ (25)
-for all $x\in{\mathbb{R}}^{Nn}$ and $t\geq 0$.
+for all $x\in{\mathbb{R}}^{Nn}$ and $t\in\mathbb{N}$.
 Hence, by introducing $\xi_{\text{avg}}:=\mathtt{col}(\tilde{x}_{\text{avg}},\tilde{z}_{\perp,_{\text{avg}}})$ and by combining the definition of $\phi_{\text{avg}}$ in (21), $\phi$ in (19), and (25), we obtain
-$\phi_{\text{avg}}(\xi_{\text{avg}})$ $=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$
+$\phi_{\text{avg}}(\xi_{\text{avg}})\!$ $=\!\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}\!-\!R^{\top}L(G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\!-\!G(\mathbf{1}x^{\star}))\end{bmatrix}$
 $\hskip 11.38092pt+\begin{bmatrix}-\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\\ -R^{\top}L\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})\end{bmatrix}.$ (26)
 We define
-$\phi_{GT}(\xi_{\text{avg}})$ $:=\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ -R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}-R^{\top}L(G(\tilde{x}+\mathbf{1}x^{\star})-G(\mathbf{1}x^{\star}))\end{bmatrix}$
+$\phi_{GT}(\xi_{\text{avg}})\!$ $:=\!\begin{bmatrix}-L\tilde{x}_{\text{avg}}-R\tilde{z}_{\perp,_{\text{avg}}}-G(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})+G(\mathbf{1}x^{\star})\\ \!-\!R^{\top}LR\tilde{z}_{\perp,_{\text{avg}}}\!-\!R^{\top}L(G(\tilde{x}_{\text{avg}}\!+\!\mathbf{1}x^{\star})\!-\!G(\mathbf{1}x^{\star}))\end{bmatrix}$
 $B\!$ $:=\!\begin{bmatrix}-I\\ -R^{\top}L\end{bmatrix},\quad u(\xi_{\text{avg}}):=\ell(\tilde{x}_{\text{avg}}+\mathbf{1}x^{\star})$
@@
 Lemma 2
-There exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $a_{1},a_{2},c_{1}>0$, and $\gamma_{0}>0$ such that, for any $\gamma\in(0,\gamma_{0})$, along the trajectories of (28) it holds
+There exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ and $a_{1},a_{2},c_{1},\gamma_{0}>0$ such that, for any $\gamma\in(0,\gamma_{0})$, along the trajectories of (28) it holds
 $a_{1}I\leq P$ $\leq a_{2}I$ (29a)
 ${\xi_{\text{avg}}^{t+1}}^{\top}P\xi_{\text{avg}}^{t+1}-{\xi_{\text{avg}}^{t}}^{\top}P\xi_{\text{avg}}^{t}$ $\leq-\gamma c_{1}\left\|\xi_{\text{avg}}^{t}\right\|^{2},$ (29b)
-for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$. $\square$
+for all $\xi_{\text{avg}}^{t}\in{\mathbb{R}}^{n_{\xi}}$. $\square$
 The proof of Lemma 2 is in Appendix -B.
 Remark 3
-Notice that Lemma 2 proves the algorithm (8) linearly converges to the minimizer of (1), since (28) is an equivalent formulation of (8). $\square$
-With this result at hand, we analyze the impact of $u(\cdot)$ thus obtaining the stability properties for the averaged system (27).
+Notice that Lemma 2 proves that algorithm (8) linearly converges to the minimizer of (1), since (28) is an equivalent formulation of (8). $\square$
+With this result at hand, we analyze the impact of $u(\cdot)$ thus obtaining the stability properties of the averaged system (27).
 Lemma 3
@@
 Then, for any $r_{\xi}>0$ and $\rho\in(0,r_{\xi})$, there exist $c_{3}\in(0,c_{1})$ and $\delta^{\star}_{1}\in(0,1]$ such that, for any $\gamma\in\min\{\gamma_{0},1\}$, $\delta\in(0,\delta^{\star}_{1})$, and $\left\|\xi_{\text{avg}}^{0}\right\|\leq r_{\xi}$, it holds
-(i) $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\sqrt{a_{2}/a_{1}}r_{\xi}}$ for all $t\geq 0$,
+(i) $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\sqrt{a_{2}/a_{1}}r_{\xi}}$ for all $t\in\mathbb{N}$,
 $\|\xi_{\text{avg}}^{t}\|\leq\sqrt{a_{2}/a_{1}}\exp\left(-t\gamma c_{3}\right)\|\xi_{\text{avg}}^{0}\|,$ (30)
@@
 10] and [41] for elaborating the proof of Theorem 1 to make clear how $\gamma$ affects the closeness of the trajectories of (18) and (27).
-Since Assumptions 12, and 3 hold, we apply Lemma 2 to claim that there exists $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $\gamma_{0}>0$, and $a_{1},a_{2},c_{1}>0$ such that, if $\gamma\in(0,\gamma_{0})$, the conditions (29) are satisfied.
+Since Assumptions 12, and 3 hold, we apply Lemma 2 to claim that there exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ and $a_{1},a_{2},c_{1},\gamma_{0}>0$ such that, if $\gamma\in(0,\gamma_{0})$, the conditions (29) are satisfied.
 Then, we evaluate the norm of the initial conditions of system (18) and (27), i.e., $\left\|\xi^{0}\right\|=\left\|\xi^{0}_{\text{avg}}\right\|$.
@@
 Then, we pick $\rho\in(0,(\bar{\rho}/2-(1+c_{2})\epsilon)/c_{2})$, $c_{1}^{\prime}\in(0,c_{1})$, and use the matrix $P$ satisfying (29) to apply Lemma 3.
-Specifically, we claim that there exist $c_{3}>0$, $\delta^{\star}_{1}\in(0,1]$, and $\gamma_{0}>0$ such that, for any $\delta\in(0,\delta^{\star}_{1})$ and $\gamma\in(0,\min\{\gamma_{0},\}]$, it holds $\xi_{\text{avg}}^{t}\in\mathcal{B}_{c_{2}r_{\xi}}$ for all $t\geq 0$ and
+Specifically, we claim that there exist $c_{3}>0$, $\delta^{\star}_{1}\in(0,1]$, and $\gamma_{0}>0$ such that, for any $\delta\in(0,\delta^{\star}_{1})$ and $\gamma\in(0,\min\{\gamma_{0},1\})$, it holds $\xi_{\text{avg}}^{t}\in\mathcal{B}_{c_{2}r_{\xi}}$ for all $t\in\mathbb{N}$ and
 $\left\|\xi_{\text{avg}}^{t}\right\|\leq c_{2}\exp(-t\gamma c_{3})\left\|\xi_{\text{avg}}^{0}\right\|,$ (31)
 for all $\xi_{\text{avg}}^{t}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$.
-Now, we proceed by finding a bound $\gamma_{1}>0$ such that, for any $\gamma\in(0,\gamma_{1})$, we guarantee the $\epsilon$-closeness between the state $\xi_{\text{avg}}^{t}$ of the averaged system (20) and the one of the original system (18), i.e., that $\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon$ holds true for all $t\geq 0$.
+Now, we proceed by finding a bound $\gamma_{1}>0$ such that, for any $\gamma\in(0,\gamma_{1})$, we guarantee the $\epsilon$-closeness between the state $\xi_{\text{avg}}^{t}$ of the averaged system (20) and the one of the original system (18), i.e., that $\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon$ holds true for all $t\in\mathbb{N}$.
 To this end, let us introduce
@@
 Then, let $r_{\xi}^{\prime}:=c_{2}r_{\xi}$ and define $\Delta:=\delta\sqrt{Nn}$.
-Under the assumption of $\xi^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\geq 0$ (later verified by a proper selection of $\gamma$), we claim that the arguments of the functions $f_{i}$ and their derivatives (embedded into the definitions of $\phi(t,\cdot)$ and $\phi_{\text{avg}}(\cdot)$ and their derivatives) lie into the compact set $\mathcal{B}_{r^{\prime}_{\xi}+\epsilon+\Delta}$.
+Under the assumption of $\xi^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in\mathbb{N}$ (later verified by a proper selection of $\gamma$), we claim that the arguments of the functions $f_{i}$ and their derivatives (embedded into the definitions of $\phi(t,\cdot)$ and $\phi_{\text{avg}}(\cdot)$ and their derivatives) lie into the compact set $\mathcal{B}_{r^{\prime}_{\xi}+\epsilon+\Delta}$.
 Thus, since the functions $f_{i}$ and its derivatives are continuous (cf.
@@
 $\left\|\frac{\partial\phi_{\text{avg}}(\xi)}{\partial\xi}\right\|,\left\|\frac{\partial\nu(t,\xi)}{\partial\xi}\right\|\bigg{\}}.$ (33)
-Consequently, for all $\xi,\xi^{\prime}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ and $t\geq 0$, it holds
+Consequently, for all $\xi,\xi^{\prime}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ and $t\in\mathbb{N}$, it holds
 $\|\upsilon(t,\xi)\|$ $\leq 2L_{\phi}\tau_{\text{per}}$ (34a)
@@
 $\theta^{\star}\geq-\tfrac{1}{c_{3}}\ln\left(\tfrac{(\bar{\rho}/2-\epsilon)/c_{2}}{c_{2}r_{\xi}}\right).$ (39)
-Let $\gamma_{2}:=\tfrac{\epsilon/3}{L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)\theta^{\star}}$, $\gamma_{3}:=\tfrac{\epsilon/3}{2L_{\phi}(1+2\tau_{\text{per}})\exp(L_{\phi}\theta^{\star})}$,$\gamma_{4}:=\tfrac{\epsilon/3}{2L_{\phi}\tau_{\text{per}}}$, $\gamma_{1}:=\min\{\gamma_{0},\gamma_{2},\gamma_{3},\gamma_{4},1\}$.
+Let $\gamma_{2}:=\tfrac{\epsilon/(3c_{2})}{L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)\theta^{\star}}$, $\gamma_{3}:=\tfrac{\epsilon/(3c_{2})}{2L_{\phi}(1+2\tau_{\text{per}})\exp(L_{\phi}\theta^{\star})}$,$\gamma_{4}:=\tfrac{\epsilon/(3c_{2})}{2L_{\phi}\tau_{\text{per}}}$, $\gamma_{1}:=\min\{\gamma_{0},\gamma_{2},\gamma_{3},\gamma_{4},1\}$.
 Pick $\gamma\in(0,\gamma_{1})$ such that $t^{\star}:=\frac{\theta^{\star}}{\gamma}\in{\mathbb{N}}$.
@@
 Then, we bound (38) as
-$\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon,$ (40)
-for all $t\in[0,t^{\star}]$.
-As a consequence, since $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}}$ for all $t\geq 0$, it holds ${\xi}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in[0,t^{\star}]$, i.e., we have verified that the bounds (34) can be used into the interval $[0,t^{\star}]$.
+$\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\frac{\epsilon}{c_{2}},$ (40)
+for all $t\in\{0,\dots,t^{\star}\}$.
+As a consequence, since $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}}$ for all $t\in\mathbb{N}$, it holds ${\xi}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in\{0,\dots,t^{\star}\}$, i.e., we have verified that the bounds (34) can be used into the interval $\{0,\dots,t^{\star}\}$.
 Moreover, the exponential law (31) and the expression of $\theta^{\star}$ (cf.
@@
 The inequality (42) guarantees that $\xi^{t^{\star}}\in\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$, hence we proved that the trajectories of (27) enters into $\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$ with linear rate.
-Next, in order to show that $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\geq t^{\star}$, we divide the set of natural numbers in intervals as $\mathbb{N}=[0,\,t^{\star}]\cup[t^{\star},2t^{\star}]\cup\dots$.
-Define $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ as the solution to (27) for $\xi_{\text{avg}}^{0}=\xi^{t^{\star}}$ and $k\in[0,t^{\star}]$.
-Thus, at the beginning of each interval $[t^{\star},2t^{\star}]$, the initial condition of (27) coincides with the one of $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ and lies into $\mathcal{B}_{\bar{\rho}/2}\subseteq\mathcal{B}_{r_{\xi}}$.
+Next, in order to show that $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\geq t^{\star}$, we divide the set of natural numbers in intervals as =$\mathbb{N}=\{0,\dots,t^{\star}\}\cup\{t^{\star},\dots,2t^{\star}\}\cup\dots$.
+Define $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ as the solution to (27) for $\xi_{\text{avg}}^{0}=\xi^{t^{\star}}$ and $k\in\{0,\dots,t^{\star}\}$.
+Thus, at the beginning of each interval $\{t^{\star},\dots,2t^{\star}\}$, the initial condition of (27) coincides with the one of $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ and lies into $\mathcal{B}_{\bar{\rho}/2}\subseteq\mathcal{B}_{r_{\xi}}$.
 Thus, we apply the same arguments above to guarantee that
-$\|\xi^{k+t^{\star}}-\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})\|\leq\epsilon,$
-for all $\gamma\in(0,\gamma^{\star})$, $k\in[0,t^{\star}]$, and $\psi_{\text{avg}}(2t^{\star},\xi^{t^{\star}})\in\mathcal{B}_{(\bar{\rho}/2-\epsilon)/c_{2}}$.
+$\|\xi^{k+t^{\star}}-\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})\|$ $\leq\epsilon$
+$\psi_{\text{avg}}(2t^{\star},\xi^{t^{\star}})$ $\in\mathcal{B}_{(\bar{\rho}/2-\epsilon)/c_{2}},$
+for all $\gamma\in(0,\gamma^{\star})$ and $k\in\{0,\dots,t^{\star}\}$.
 By using Lemma 3, we guarantee that system (27) cannot escape from $\mathcal{B}_{\bar{\rho}/2-\epsilon}$, namely $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\bar{\rho}/2-\epsilon}$ for all $t\geq t^{\star}$.
-Thus, we get $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\in[t^{\star},2t^{\star}]$.
-By recursively applying the same arguments above for each interval $[jt^{\star},(j+1)t^{\star}]$ with $j=2,3,\dots$ and using $\left\|x_{i}^{t}-x^{\star}\right\|\leq\left\|\xi^{t}\right\|$ for all $i\in\{1,\ldots,N\}$ and $t\geq 0$, we get
+Thus, we get $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\in\{t^{\star},\dots,2t^{\star}\}$.
+By recursively applying the same arguments above for each interval $\{jt^{\star},\dots,(j+1)t^{\star}\}$ with $j=2,3,\dots$ and using $\left\|x_{i}^{t}-x^{\star}\right\|\leq\left\|\xi^{t}\right\|$ for all $i\in\{1,\ldots,N\}$ and $t\in\mathbb{N}$, we get
 $\left\|x_{i}^{t}-x^{\star}\right\|\leq\bar{\rho}/2,$ (43)
-for all $i\in\{1,\ldots,N\}$ and $t\geq 0$. The change of coordinates (6), (43), and the triangle inequality lead to
-$\left\|w_{i}^{t}-x^{\star}\right\|$ $\leq\frac{\bar{\rho}}{2}+\delta\left\|d_{i}^{t}\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}\bar{\rho}/2+\delta\sqrt{n},$ (44)
+for all $i\in\{1,\ldots,N\}$ and $t\in\mathbb{N}$. The change of coordinates (6), (43), and the triangle inequality lead to
+$\left\|w_{i}^{t}-x^{\star}\right\|$ $\leq\frac{\bar{\rho}}{2}+\delta\left\|d_{i}^{t}\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\bar{\rho}}{2}+\delta\sqrt{n},$ (44)
 where in $(a)$ we use the boundedness of the dither signals. The proof follows from (44) by setting $\delta^{\star}:=\min\left\{\delta^{\star}_{1},\frac{\bar{\rho}}{2\sqrt{n}}\right\}$.
@@
 In these scenarios, the user discomfort term is usually not known in advance but can be only accessed by measurements.
-Specifically, we associate to each agent $i\in\{1,\ldots,N\}$ a cost function in the form $f_{i}(w)=w^{\top}Q_{i}w+r_{i}^{\top}w+\log(\sum_{\ell=1}^{n}a_{i\ell}e^{b_{i\ell}w_{\ell}})$ with $Q_{i}=Q_{i}^{\top}\in^{n\times n}$, $r_{i}\in^{n}$ and $a_{i\ell},b_{i\ell}$, for all $\ell\in\{1,\ldots,n\}$.
+Specifically, we associate to each agent $i\in\{1,\ldots,N\}$ a cost function in the form $f_{i}(w)=w^{\top}Q_{i}w+r_{i}^{\top}w+\log(\sum_{\ell=1}^{n}a_{i\ell}e^{b_{i\ell}w_{\ell}})$ with $Q_{i}=Q_{i}^{\top}\in^{n\times n}$, $r_{i}\in^{n}$ and $a_{i\ell},b_{i\ell}>0$, for all $\ell\in\{1,\ldots,n\}$.
 In the following, we provide different sets of simulations to study in detail the features of Extremum Seeking Tracking.
@@
 We consider different optimization variable sizes, namely $n=10,20$.
-Part of these performances has been run on the Marconi100 HPC Cluster of the Italian Cineca.
+Part of these simulations has been run on the Marconi100 HPC Cluster of the Italian Cineca.
 We used $10$ nodes of the cluster and, for each node, we used $25$ cores and $4$ GPUs.
@@
 2.
-In detail, increasing the number of agents increases the Lipschitz constant of the system to be averaged.
+In detail, as the number of agents increases, the Lipschitz constant of the system to be averaged increases too.
 Moreover, a larger domain of initial conditions also implies a potentially larger $L_{\phi}$ constant (cf.
@@
 For each value of $N$, we generate communication graphs with a diameter $d$ such that the ratio $N/d$ is constant while varying $N$.
-Results are depicted in Fig.
+The results depicted in Fig.
 3 show that the algorithm slows down as the number of agents increases.
-Figure 3: Monte Carlo simulations for varying number of agents: mean and $1$-standard deviation band of cost error (left) and variable error (right).
+Figure 3: Monte Carlo simulations for a varying number of agents: mean and $1$-standard deviation band of cost error (left) and variable error (right).
 V-C Monte Carlo simulations varying the size $n$
@@
 Figure 4: Monte Carlo simulations varying $n$: mean and $1$-standard deviation band of cost error (left) and variable error (right).
-We conclude this part by providing the results of a single trial in the case with $n=1$ to show the evolution of the solution estimate of each agent in error coordinate, with respect to the optimal solution to the problem.
+We conclude this part by providing in Fig.
+5 the results of a single trial in the case with $n=1$ to show the evolution of the solution estimate of each agent in error coordinate with respect to the optimal solution to the problem.
 Figure 5: Evolution of the agents’ estimates $w_{i}^{t}$ in coordinate error with respect to the optimal solution $x^{\star}$.
@@
 7 shows that, although the algorithm by [19] exhibits a faster convergence in the beginning phase of the simulations, our scheme has a better convergence rate and final accuracy.
-Indeed, although both Extremum Seeking Tracking and 1P-DSGT require a single function query for each iteration, we recall that 1P-DSGT adopts a stochastic dither whose theoretical gradient estimation properties are guaranteed only asymptotically and in the case of ergodic processes.
-On the opposite, our periodic dither guarantees that the theoretical gradient estimation is (practically) achieved with $\tau_{\text{per}}$ functions’ samples (in an averaged sense).
+Both 1P-DSGT and Extremum Seeking Tracking employ a single function query per agent to approximate the gradient.
+These methods rely on the concept that averaging across iterations around the “quasi-static” local solution estimate $x_{i}^{t}$ provides an accurate gradient approximation.
+However, in the context of their theoretical frameworks, 1P-DSGT attains this approximation using an infinite number of samples (mean of an ergodic process), whereas Extremum Seeking Tracking only needs $\tau_{\text{per}}$ samples.
 This makes the convergence rate of our algorithm faster.
@@
 Finally, given a compact set $\mathcal{S}_{i}\subset{\mathbb{R}}^{n}$, let us bound $\left\|\ell_{i}(x_{i})\right\|$ for all $x_{i}\in\mathcal{S}_{i}$.
-Note that $\left\|\delta d_{i}^{t}\right\|\leq\delta\sqrt{n}$ for all $t\geq 0$ and let $\mathcal{S}^{\prime}_{i}\subset{\mathbb{R}}^{n}$ be a compact set such that (i) $\mathcal{S}_{i}\subseteq\mathcal{S}_{i}^{\prime}\subset{\mathbb{R}}^{n}$, and (ii) $x_{i}+\delta d_{i}^{t}\in\mathcal{S}^{\prime}$ for all $x_{i}\in\mathcal{S}_{i}$, $\delta\in(0,1]$, and $t\geq 0$.
+Note that $\left\|\delta d_{i}^{t}\right\|\leq\delta\sqrt{n}$ for all $t\in\mathbb{N}$ and let $\mathcal{S}^{\prime}_{i}\subset{\mathbb{R}}^{n}$ be a compact set such that (i) $\mathcal{S}_{i}\subseteq\mathcal{S}_{i}^{\prime}\subset{\mathbb{R}}^{n}$, and (ii) $x_{i}+\delta d_{i}^{t}\in\mathcal{S}^{\prime}$ for all $x_{i}\in\mathcal{S}_{i}$, $\delta\in(0,1]$, and $t\in\mathbb{N}$.
 Thus, we can write
@@
 We underline that, since the set $\mathcal{S}_{i}^{\prime}$ is compact and $f_{i}$ is smooth, $L_{i,\mathcal{S}_{i}}^{\prime}$ exists and is finite.
-The bound of $\ell_{i}(x_{i})$ follows by defining $L_{i,\mathcal{S}_{i}}:=2L_{i,\mathcal{S}_{i}}^{\prime}\sqrt{n}$ and combining the result (48) with the bound about the norm of the dither signal, i.e., $\left\|d_{i}^{t}\right\|\leq\sqrt{n}$ for all $t\geq 0$.
+The bound of $\ell_{i}(x_{i})$ follows by defining $L_{i,\mathcal{S}_{i}}:=2L_{i,\mathcal{S}_{i}}^{\prime}\sqrt{n}$ and combining the result (48) with the bound about the norm of the dither signal, i.e., $\left\|d_{i}^{t}\right\|\leq\sqrt{n}$ for all $t\in\mathbb{N}$.
 -B Proof of Lemma 2
@@
 In detail, [38, Th.
-3.1] introduces $\bar{T}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ to define $\bar{\xi}:=\bar{T}\xi$, and a matrix $\bar{P}=\bar{P}^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ such that
+3.1] introduces a full-rank matrix $\bar{T}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ to define $\bar{\xi}:=\bar{T}\xi$, and a matrix $\bar{P}=\bar{P}^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ such that
 $\bar{a}_{1}I$ $\leq\bar{P}\leq\bar{a}_{2}I$ (49a)
 $2\bar{\xi}_{\text{avg}}^{\top}\bar{P}\bar{T}\phi_{\text{GT}}(\bar{T}^{-1}\bar{\xi}_{\text{avg}})$ $\leq-\bar{a}_{3}\left\|\bar{\xi}_{\text{avg}}\right\|^{2},$ (49b)
-for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$ and some $\bar{a}_{1},\bar{a}_{2},\bar{a}_{3}>0$. Let $P:=\bar{T}^{-1}\bar{P}T$. Then, there exist $a_{1},a_{2},a_{3}>0$ such that
+for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$ and some $\bar{a}_{1},\bar{a}_{2},\bar{a}_{3}>0$. Let $P:=\bar{T}^{\top}\bar{P}T$. Then, there exist $a_{1},a_{2},a_{3}>0$ such that
 $a_{1}I$ $\leq P\leq a_{2}I$ (50a)
@@
 Hence, by combining (55) and (56), if $\delta\in(0,\delta^{\star}_{1})$, then, for all $\xi_{\text{avg}}^{t}\in\Omega_{r^{\prime}_{\xi_{\text{avg}}}}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$, it holds
-$\Delta V(\xi_{\text{avg}}^{t})<-\gamma c_{1}^{\prime}\left\|\xi_{\text{avg}}^{t}\right\|^{2}.$ (57)
-Thus, the inequality (57) ensures that the set $\Omega_{r_{\xi}}$ is invariant for system (27).
-Hence, if we pick $\xi_{\text{avg}}^{0}\in\mathcal{B}_{r_{\xi}}$, we prove that $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$ for all $t\geq 0$.
-Consequently, the bound (54) holds for all $t\geq 0$ and, in turn, also the inequality (57) is verified for all $t\geq 0$, namely we proved that the trajectories of system (27) enter the ball $\mathcal{B}_{\rho}$ exponentially fast.
+$\Delta V(\xi_{\text{avg}}^{t})\leq-\gamma c_{1}^{\prime}\left\|\xi_{\text{avg}}^{t}\right\|^{2}.$ (57)
+Thus, the inequality (57) ensures that the set $\Omega_{r_{\xi}}$ is forward-invariant for system (27).
+Hence, if we pick $\xi_{\text{avg}}^{0}\in\mathcal{B}_{r_{\xi}}$, we prove that $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$ for all $t\in\mathbb{N}$.
+Consequently, the bound (54) holds for all $t\in\mathbb{N}$ and, in turn, also the inequality (57) is verified for all $t\in\mathbb{N}$, namely we proved that the trajectories of system (27) enter the ball $\mathcal{B}_{\rho}$ exponentially fast.
 The result (30) follows from the inequality (57) and (29a) by setting $c_{3}:=c_{1}^{\prime}/(2a_{2})$.